### v2

# Representing Sugihara monoids

Abstract
We show that all Sugihara monoids can be represented as algebras of binary relations, with the monoid operation given by relational composition. Moreover, the binary relations are weakening relations. The first step is to obtain an explicit relational representation of all finite odd Sugihara chains. Our construction mimics that of Maddux (2010), where a relational representation of the finite even Sugihara chains is given. We define the class of representable Sugihara monoids as those which can be represented as so-called direct reducts of distributive involutive FL-algebras of binary relations. We then show that the class of representable distributive involutive FL-algebras is closed under ultraproducts. This fact is used to demonstrate that the two infinite Sugihara monoids that generate the quasivariety are also representable. From this it follows that all Sugihara monoids are representable.
1 Introduction
Sugihara monoids and Sugihara algebras have found many applications in the study of relevance logics. The variety of Sugihara algebras is generated by $\mathbf{Z}=\langle\mathbb{Z},\wedge,\vee,\cdot,\to,{\sim}\rangle$ and provides complete algebraic semantics for the logic $\mathbf{R}$-mingle (often denoted $\mathbf{RM}$[1, 4]. Sugihara monoids can be described as commutative distributive idempotent residuated lattices with an order-reversing involution. That is, a truth constant, which is the identity element of the monoid, is included in the signature. They are known to give algebraic semantics for $\mathbf{RM^{t}}$, the logic $\mathbf{R}$-mingle with added Ackermann constants [1].
Odd Sugihara monoids are those for which the monoid identity (truth constant) is fixed by the involution. They were studied by Galatos and Raftery [12] where they showed a categorical equivalence with relative Stone algebras. An application provided results about Beth definability and interpolation for the uninorm-based logic $\mathbf{IUML}$. Dual topological representations of Sugihara monoids were studied extensively by Fussner and Galatos [7] (see also Fussner’s PhD thesis [6]). Dual spaces were constructed in two ways—making use of of both Priestley spaces and Esakia spaces.
Our goal in this paper is to show that it is possible to find a representation of any Sugihara monoid, with the monoid operation given by relational composition. We do this by first giving an explicit relational representation for each finite odd Sugihara chain. Given a poset $\langle X,\leqslant\rangle$, a binary relation $R$ on $X$ is said to be a weakening relation if ${\leqslant}\circ R\circ{\leqslant}=R$. The elements of our algebra will be weakening relations on a poset which comes from a twisted product of a poset with its dual. Our construction essentially combines the method of Maddux [16] with recent work by the current authors [3]. There we were able to represent so-called distributive quasi relation algebras (DqRAs) where the elements of the algebra are binary relations and the monoid operation is given by relational composition. Quasi relation algebras (qRAs) were introduced by Galatos and Jipsen [8] as a generalisation of relation algebras. A major feature of these algebras is that the variety of qRAs has a decidable equational theory (whereas the variety of relation algebras does not).
In 2010, Maddux [16, Theorem 6.2] gave a relational representation of the finite even Sugihara chains $\mathbf{S}_{2n+2}$ ($n\in\mathbb{Z}^{+}$). We outline that representation here in very general terms, to help us describe the modification we make in order to represent the finite odd Sugihara chains. For $n\in\mathbb{Z}^{+}$, consider the set $\mathbb{Q}^{n}$ with the lexicographic ordering, denoted $\leqslant_{n}$. Now $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$ is a chain. With $\mathbf{X}=\langle X,\leqslant\rangle=\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$, consider the usual product poset $\mathbf{X}^{\partial}\times\mathbf{X}$ where $\mathbf{X}^{\partial}=\langle X,\geqslant\rangle$. It is then possible to embed $\mathbf{S}_{2n+2}$ in the set of upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$ in such a way that the elements of the algebra are binary relations and the monoid operation is given by relational composition. This construction, which is just a re-casting of [16], is described in more detail in Example 2.11.
For the odd Sugihara chain $\mathbf{S}_{2n+3}$ (where $n\in\mathbb{Z}^{+}$), we will take two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$, with a map $\alpha$ swapping the occurrence of each element from one copy to the other. This map $\alpha$ is an order isomorphism of the poset consisting of two chains. The algebra $\mathbf{S}_{2n+3}$ is then embedded in the upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$, where this time $\mathbf{X}$ consists of two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$. The key step is that $\alpha$ is used to define the order-reversing involution in a way that ensures the algebra will be an odd Sugihara monoid.
Indeed, it is the usage of such a map $\alpha$ that distinguishes our construction from other relational representations of similar algebras. As mentioned above, it provides a method for having algebras with ${\sim}1=1$. In the case of representable weakening relation algebras [9, 10, 13], the complement-converse is used as the involutive order-reversing operation on binary relations. It is not difficult to show that no binary relation can be fixed by complement-converse (cf. [16, p.46] or  [5, p.355]). We note that in the relational representation of quasi-Boolean algebras by Dunn [5], that construction solves the problem of being able to have an order-reversing involution that fixes certain relations. It does so by taking the complement inside a symmetric relation, which does not have to be $X^{2}$.
The map $\alpha$ also plays a crucial role in the representation of distributive involutive FL-algebras. These algebras have two order-reversing maps $\sim$ and -- (see Sections 2.1 and 2.4). If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic. The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7). Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
[…]
Our paper is laid out as follows. In Section 2 we give some basic definitions and background on Sugihara monoids and binary relations. We also outline how we use partially ordered equivalence relations (equipped with an order automorphism $\alpha$) to obtain relational representations of distributive involutive FL-algebras, and give some examples. In Section 3 we build the finite odd Sugihara chains $\mathbf{S}_{2n+1}$ ($n\in\mathbb{Z}^{+}$) as algebras of binary relations. The proofs contained in Section 3 are rather technical, but this is mostly due to the large number of cases that need to be considered. The feature of our construction is the relative simplicity of the relations from which the algebra is built.
In the final section, we first give a definition of a representable Sugihara monoid. We then show that the class of representable DInFL-algebras (see Definition 2.14) is closed under ultraproducts (Theorem 4.11). Then, combining the main result of Section 3 with a representation of finite even Sugihara chains by Maddux [16], we are able to show that all Sugihara monoids are representable using the partially ordered equivalence relations described in Section 2.3.
2 Preliminaries
Before we provide the details of our construction, we give some background on residuated lattices, Sugihara monoids, the algebra of binary relations, some necessary results from our previous work [3], as well as a definition of representability of a DInFL-algebra.
2.1 Sugihara monoids and DInFL-algebras
A residuated lattice is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,\backslash,/\rangle$ such that $\langle A,\wedge,\vee\rangle$ is a lattice, $\langle A,\cdot,1\rangle$ is a monoid and the following holds for all $a,b,c\in A$:
$a\cdot b\leqslant c\quad\Longleftrightarrow\quad a\leqslant c/b\quad\Longleftrightarrow\quad b\leqslant a\backslash c.$
[…]
$a_{i}\cdot a_{j}=\begin{cases}a_{i}&\text{if }\,|j|<|i|\\ a_{j}&\text{if }\,|i|<|j|\\ a_{\text{min}\{i,j\}}&\text{if }\,|j|=|i|.\end{cases}$
Lastly, the implication $\to$ is defined by
$a_{i}\to a_{j}=\begin{cases}{\sim}a_{i}\vee a_{j}&\text{if }i\leqslant j\\ {\sim}a_{i}\wedge a_{j}&\text{if }i>j.\end{cases}$
If $n$ is odd then the monoid identity is given by $1=a_{0}$, and if $n$ is even, then the monoid identity is $1=a_{1}$. This gives us algebras $\mathbf{S}_{n}=\langle S_{n},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$. (We remark that when we talk of a Sugihara chain, we mean an algebra that is a Sugihara monoid and is linearly ordered. This is different to the terminology used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
[…]
2.2 Binary relations
Below we recall some basic definitions and facts about binary relations and the various operations defined on them. We will use the following special binary relations on a set $X$: the empty relation $\varnothing$, the identity relation (diagonal relation) $\mathrm{id}_{X}=\left\{\left(x,x\right)\mid x\in X\right\}$, and the universal relation $X^{2}$. The complement of $R$ will be denoted by $R^{c}$ and the converse of $R$ is $R^{\smile}=\left\{\left(x,y\right)\mid\left(y,x\right)\in R\right\}$. The composition of $R,S\subseteq X^{2}$ is $R\circ S=\left\{\left(x,y\right)\mid\left(\exists z\in X\right)\left(\left(x,z\right)\in R\textnormal{ and }\left(z,y\right)\in S\right)\right\}$. Besides composition, binary relations have the familiar operations of union and intersection that will be used to model join and meet in the relational representations.
Proposition 2.1
Let $R,S,T\subseteq X^{2}$. Then the following hold:
  1. (i)
    $R^{\smile\smile}=R$
  2. (ii)
    $R^{\smile c}=R^{c\smile}$
  3. (iii)
    $\left(R\cup S\right)^{\smile}=R^{\smile}\cup S^{\smile}$
  4. (iv)
    $\left(R\cap S\right)^{\smile}=R^{\smile}\cap S^{\smile}$
  5. (v)
    $\mathrm{id}_{X}\circ R=R\circ\mathrm{id}_{X}=R$
  6. (vi)
    $\left(R\circ S\right)\circ T=R\circ\left(S\circ T\right)$
  7. (vii)
    $\left(R\circ S\right)^{\smile}=S^{\smile}\circ R^{\smile}$
  8. (viii)
    $(R\cup S)\circ T=(R\circ T)\cup(S\circ T)$
  9. (ix)
    $R\circ(S\cup T)=(R\circ S)\cup(R\circ T)$
A crucial part of our construction is the use of an additional bijective function $\alpha:X\to X$. We will use $\alpha$ to denote both the function acting on elements of $X$ and the binary relation on $X$. The lemma below applies in particular when $\gamma$ is a bijective function from $X$ to $X$.
Lemma 2.2 ([3, Lemma 3.2])
Let $X$ be a set and $R,\gamma\subseteq X^{2}$. If $\gamma$ satisfies $\gamma^{\smile}\circ\gamma=\mathrm{id}_{X}$ and $\gamma\circ\gamma^{\smile}=\mathrm{id}_{X}$ then the following hold:
  1. (i)
    $\left(\gamma\circ R\right)^{c}=\gamma\circ R^{c}$
  2. (ii)
    $\left(R\circ\gamma\right)^{c}=R^{c}\circ\gamma$
2.3 Constructing InFL-algebras of relations
[…]
The definitions and results below all involve an equivalence relation $E$ on a set $X$. The special case of $E=X^{2}$ is vastly simpler and, fortunately, this is the setting used in Section 3. However, we need the more general setting for the representability results in Section 4.
Let $\mathbf{X}=\left\langle X,\leqslant\right\rangle$ be a poset with an equivalence relation $E$ satisfying ${\leqslant}\subseteq E$. The set $E$ can be partially ordered for any $(u,v),(x,y)\in E$, as follows:
$(u,v)\preccurlyeq(x,y)\qquad\textnormal{iff}\qquad x\leqslant u\textnormal{ and }v\leqslant y.$
Then $\mathbf{E}=\left\langle E,\preccurlyeq\right\rangle$ is a poset and the set of its upsets, denoted $\mathsf{Up}\left(\mathbf{E}\right)$, ordered by inclusion, is a distributive lattice. Notice that when $E=X^{2}$ then $\langle X^{2},\preccurlyeq\rangle$ is exactly the product poset $\mathbf{X}^{\partial}\times\mathbf{X}$ where the order on $\mathbf{X}^{\partial}$ is $\geqslant$.
[…]
The lemma below is stated in the original source [3] with an additional map $\beta:X\to X$ which is a dual order automorphism. This is not required for the setting of DInFL-algebras. For a poset $\mathbf{Y}=\langle Y,\leqslant_{Y}\rangle$, we denote by $\mathsf{Down}(\mathbf{Y})$ the set of downsets of $\mathbf{Y}$. The lemma shows how upsets/downsets of $\mathbf{E}=\langle E,\preccurlyeq\rangle$ are produced when composition, converse and complement are applied to upsets/downsets of $\mathbf{E}=\langle E,\preccurlyeq\rangle$. It also details what happens when the graph of an order automorphism is composed with an upset. For $R\subseteq X^{2}$, we will abuse notation and write $R^{c}$ for the complement of $R$ in $E$.
Lemma 2.3 ([3, Lemma 3.5])
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset, $E$ an equivalence relation with ${\leqslant}\subseteq E$ and let $\alpha:X\rightarrow X$ be an order automorphism of $\mathbf{X}$ with $\alpha\subseteq E$.
  1. (i)
    If $R,S\in\mathsf{Up}\left(\mathbf{E}\right)$, then $R\circ S\in\mathsf{Up}\left(\mathbf{E}\right)$.
  2. (ii)
    $R\in\mathsf{Up}\left(\mathbf{E}\right)$ iff $R^{c}\in\mathsf{Down}\left(\mathbf{E}\right)$.
  3. (iii)
    $R\in\mathsf{Up}\left(\mathbf{E}\right)$ iff $R^{\smile}\in\mathsf{Down}\left(\mathbf{E}\right)$.
  4. (iv)
    If $R\in\mathsf{Up}\left(\mathbf{E}\right)$, then $\alpha\circ R\in\mathsf{Up}\left(\mathbf{E}\right)$ and $R\circ\alpha\in\mathsf{Up}\left(\mathbf{E}\right)$.
The next lemma shows not only that $\leqslant$ is an upset of $\mathbf{E}$, but also that it is the identity of composition in $\mathsf{Up}(\mathbf{E})$.
Lemma 2.4 ([3, Lemma 3.6])
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset, $E$ and equivalence relation with ${\leqslant}\subseteq E$ and $\alpha$ an order automorphism on $X$ with $\alpha\subseteq E$. Then
  1. (i)
    $\leqslant\;\in\mathsf{Up}\left(\mathbf{E}\right)$;
  2. (ii)
    $\leqslant\circ\,R=R\;\circ\leqslant\;=R$ for all $R\in\mathsf{Up}\left(\mathbf{E}\right)$.
[…]
The next lemma shows that if $\alpha$ is an order automorphism of a poset $\langle X,\leqslant\rangle$ with an equivalence relation $E$ such that ${\leqslant}\subseteq E$, then $\alpha$ can be used to define an element of $\mathsf{Up}(\mathbf{E})$ to act as a 00 that will give rise to an InFL-algebra.
Lemma 2.5 ([3, Lemma 3.9])
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with an equivalence relation $E$ on $X$ satisying ${\leqslant}\subseteq E$. Further, let $\alpha:X\rightarrow X$ be an order automorphism of $\mathbf{X}$ with $\alpha\subseteq E$. Then $\alpha\,\circ\leqslant^{c\smile}=$ $\leqslant^{c\smile}\circ\,\alpha\in\mathsf{Up}\left(\mathbf{E}\right)$.
We will now add $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ to the signature of the distributive residuated lattice $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,\backslash,/\right\rangle$ to obtain the FL-algebra $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,0,\backslash,/\right\rangle$. The next lemma gives an alternative way of calculating ${\sim}R$ and $-R$ if we set $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$.
Lemma 2.6 ([3, Lemma 3.10])
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with $E$ an equivalence relation and $\alpha:X\rightarrow X$ an order automorphism of $\mathbf{X}$ where ${\leqslant}\subseteq E$ and $\alpha\subseteq E$. If $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ then ${\sim}R=R^{c\smile}\circ\alpha$ and $-R=\alpha\circ R^{c\smile}$ for all $R\in\mathsf{Up}\left(\langle X^{2},\preccurlyeq\rangle\right)$.
[…]
Theorem 2.7 ([3, Theorem 3.12])
Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with $E$ an equivalence relation on $X$ such that ${\leqslant}\subseteq E$. Further, let $\alpha:X\rightarrow X$ be an order automorphism of $\mathbf{X}$ such that $\alpha\subseteq E$.
  1. (i)
    If  $0={\alpha}\,\circ\,{\leqslant^{c\smile}}={\leqslant^{c\smile}}\circ{\alpha}$ then $\mathfrak{D}(\mathbf{E})=\langle\mathsf{Up}(\mathbf{E}),\cap,\cup,\circ,\leqslant,{\sim},-\rangle$ is a distributive InFL-algebra.
  2. (ii)
    The algebra in (i) is cyclic if and only if $\alpha=\mathrm{id}_{X}$.
We define the class of equivalence distributive involutive FL-algebras, denoted $\mathsf{EDInFL}$, to be those algebras of the form $\mathfrak{D}(\mathbf{E})$ from part (i) of Theorem 2.7. If $E=X^{2}$ then we say that $\mathfrak{D}(\mathbf{E})$ is a full distributive involutive FL-algebra, and denote the class of such algebras by $\mathsf{FDInFL}$.
Before we give some examples, we need the following definition. It is an adaptation of [14, Definition 8].
Definition 2.8
Let $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,{\sim},-\rangle$ be a DInFL-algebra. We call $\mathbf{A}_{r}=\langle A,\wedge,\vee,\cdot,\to,1,{\sim}\rangle$ the direct reduct of $\mathbf{A}$, where $\to$ is defined for $a,b\in A$ by $a\to b={\sim}(-b\cdot a)$.
If a DInFL-algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,{\sim},-\rangle$ is commutative, then it is cyclic. Hence it is easy to see that if $\mathbf{A}$ is commutative and idempotent, then the direct reduct of $\mathbf{A}$ is a Sugihara monoid. If in addition, $\mathbf{A}$ is a chain, then $a\to b={\sim}a\vee b$ when $a\leqslant b$ and $a\to b={\sim}a\wedge b$ when $a>b$.
In the concrete case, the operation $\Rightarrow$ can be added to $\mathfrak{D}(\mathbf{E})$ using the following standard definition of the left residual of composition: for $R,S\in\mathsf{Up}(\mathbf{E})$ let $R\Rightarrow S=\left(R^{\smile}\circ S^{c}\right)^{c}$. Using Lemma 2.6 it is easy to see that this coincides with the expression given in Definition 2.8.
Example 2.9
Consider the two-element Sugihara chain $\mathbf{S}_{2}$. The underlying set is given by $S_{2}=\{a_{-1},a_{1}\}$ where $a_{1}$ is the identity of the monoid operation. Consider the one-element poset with $X=\{x\}$ and order ${\leqslant}=\mathrm{id}_{X}$. Let $E=X^{2}$ and $\alpha=\mathrm{id}_{X}$. Then $\left\langle\mathsf{Up}\left(\left\langle X^{2},\leqslant\right\rangle\right),\cap,\cup,\circ,\leqslant,\sim,-\right\rangle$ is a two-element distributive InFL-algebra whose direct reduct is isomorphic to $\mathbf{S}_{2}$.
Using Theorem 2.7, we obtain a relational representation of the odd Sugihara chain $\mathbf{S}_{3}$. We observe that this representation is not covered by the construction given in Section 3.
Example 2.10
[…]
The next example is based on the representation of the finite even Sugihara chains $\mathbf{S}_{2n+2}$ ($n\in\mathbb{Z}^{+}$) by Maddux [16, Theorem 6.2]. The exposition below shows how that representation can be re-cast into our setting. We will need this example in Section 4.2.
Example 2.11
Let $n\in\mathbb{Z}^{+}$ and consider $\mathbb{Q}^{n}=\left\{\left(q_{1},\ldots,q_{n}\right)\mid q_{1},\ldots,q_{n}\in\mathbb{Q}\right\}$. We will denote the $n$-tuple $\left(q_{1},\dots,q_{n}\right)$ by $q$. Now define binary relations $\mathrm{id}_{\mathbb{Q}^{n}}$ and $L_{1}$ on $\mathbb{Q}^{n}$ as follows:
[…]
For $i\in\{2,\ldots,n\}$, the relation $L_{i}$ on $\mathbb{Q}^{n}$ is defined follows:
$(p,q)\in L_{i}\qquad\textnormal{ iff }\qquad\left(p_{1},\ldots,p_{i-1}\right)=\left(q_{1},\ldots,q_{i-1}\right)\textnormal{ and }p_{i}<q_{i}$
The relations in $\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$ are pairwise disjoint. We set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$. Let $\mathbf{X}=\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$, $E=\left(\mathbb{Q}^{n}\right)^{2}$ and $\alpha=\mathrm{id}_{\mathbb{Q}^{n}}$. Notice that with $\alpha=\mathrm{id}_{\mathbb{Q}^{n}}$ we get ${\sim}\leqslant_{n}=\left(\leqslant_{n}\right)^{c\smile}={<_{n}}$. Moreover, by Theorem 2.7, the algebra $\mathfrak{D}(\mathbf{E})$ is a cyclic DInFL-algebra.
For $i\in\left\{-n-1,\ldots,-1,1,\ldots,n+1\right\}$, define
$T_{i}=\begin{cases}\varnothing&\text{if }\,i=-n-1\\ \displaystyle\bigcup_{j=1}^{n+1+i}L_{j}&\text{if }\,-n\leqslant i\leqslant-1\\ \leqslant_{n}&\text{if }\,i=1\\ \left(T_{-i}\right)^{c\smile}&\text{if }\,2\leqslant i\leqslant n+1\end{cases}$
These relations are upsets of $\langle E,\preccurlyeq\rangle$ and form a Sugihara chain of length $22n+$ 2, i.e.,
$\varnothing=T_{-n-1}\subseteq T_{-n}\subseteq\cdots\subseteq T_{-1}\subseteq T_{1}\subseteq\cdots\subseteq T_{n}\subseteq T_{n+1}=\left(\mathbb{Q}^{n}\right)^{2}.$
More precisely, the even Sugihara chain $\mathbf{S}_{2n+2}$ can be embedded into the direct reduct of the algebra $\mathfrak{D}(\mathbf{E})$ by setting $\varphi\left(a_{i}\right)=T_{i}$ for each $i\in\left\{-n-1,\ldots,-1,1,\ldots,n+1\right\}$.
2.4 Representable DInFL-algebras
Here we recall definitions and results from [3, Section 4]. These will be needed in Section 4. As before, the major difference is that the results there were proved for distributive quasi relation algebras, but here we restrict to the simpler case of distributive InFL-algebras.
Suppose we have an index set $I$ and for each $i\in I$, let $\mathbf{X}_{i}=\langle X_{i},\leqslant_{i}\rangle$ be a poset and $E_{i}$ an equivalence relation with $\leqslant_{i}\;\subseteq E_{i}$. We further require that $X_{i}\cap X_{j}=\varnothing$ whenever $i\neq j$. Then set
$X:=\displaystyle\bigcup_{i\in I}X_{i},\quad\displaystyle\leqslant\;:=\bigcup_{i\in I}\leqslant_{i}\quad\textnormal{and}\quad E:=\displaystyle\bigcup_{i\in I}E_{i}.$
If we have maps $\gamma_{i}:X_{i}\rightarrow X_{i}$ ($i\in I$) define $\gamma:X\to X$ by setting, for each $x\in X$, $\gamma(x)=\gamma_{i}(x)$ whenever $x\in X_{i}$. Clearly $\gamma=\displaystyle\bigcup_{i\in I}\gamma_{i}$. These definitions will be used throughout the theorems and their proofs in this section.
Theorem 2.12
[3, Theorem 4.1] Let $I$ be an index set. For each $i\in I$, let $\mathbf{X}_{i}=\langle X_{i},\leqslant_{i}\rangle$ be a poset and $E_{i}$ an equivalence relation such that $\leqslant_{i}\;\subseteq E_{i}$. Assume $X_{i}\cap X_{j}=\varnothing$ for all $i\neq j$. For each $i\in I$ we have $\alpha_{i}:X_{i}\rightarrow X_{i}$ an order automorphism of $\mathbf{X}_{i}$. Then:
  1. (i)
    $\mathbf{X}=\langle X,\leqslant\rangle$ is a poset and $E$ is an equivalence relation on $X$ such that $\leqslant\;\subseteq E$;
  2. (ii)
    $\alpha$ is an order automorphism of $\mathbf{X}$ such that $\alpha\subseteq E$;
  3. (iii)
    for the DInFL-algebras $\mathfrak{D}\left(\mathbf{E}\right)$ and $\prod_{i\in I}\mathfrak{D}\left(\mathbf{E}_{i}\right)$, we have $\mathfrak{D}\left(\mathbf{E}\right)\cong\displaystyle\prod_{i\in I}\mathfrak{D}\left(\mathbf{E}_{i}\right)$.
The following theorem relies on part (iii) of Theorem 2.12 and is used to representable DInFL-algebras.
Theorem 2.13
[3, Theorem 4.4] $\mathbb{P}\left(\mathsf{FDInFL}\right)=\mathsf{EDInFL}$
The definition below is a generalisation of [3, Definition 4.5], given there for DqRAs.
Definition 2.14
A DInFL-algebra $\mathbf{A}=\left\langle A,\wedge,\vee,\cdot,1,{\sim},{-}\right\rangle$ is representable if $\mathbf{A}\in\mathbb{ISP}\left(\mathsf{FDInFL}\right)$ or, equivalently, $\mathbf{A}\in\mathbb{IS}\left(\mathsf{EDInFL}\right)$.
In Theorem 4.11 we will show that the class of representable DInFL-algebras is closed under ultraproducts. That will be a key step in showing that all Sugihara monoids can be represented as algebras of binary relations.
We end this section by observing that if one considers posets $\langle X,=\rangle$ (i.e. with the discrete order), and $\alpha=\mathrm{id}_{X}$, then the constructions and definitions above, with some slight modification, will coincide with those for relation algebras. The modifications include using converse, complement, and bounds in the signature.
3 Constructing finite odd Sugihara chains
Let $\langle C,\leqslant\rangle$ be a chain. Set $X:=\left(C\times\left\{-1\right\}\right)\cup\left(C\times\left\{1\right\}\right)$. If $b\in\left\{-1,1\right\}$, we will write $x^{b}$ instead of $\left(x,b\right)$. Define a binary relation $\leqslant_{X}$ on $X$ by
$x^{b}\leqslant_{X}y^{d}\quad\text{ iff }\quad x^{b}=y^{d}\quad\textnormal{ or }\quad x<y$
It is not difficult to see that $\langle X,\leqslant_{X}\rangle$ is a poset. We can define a map $\alpha:X\to X$ by setting, for each $x^{b}\in X$, $\alpha(x^{b})=x^{-b}$. The map $\alpha$ will be a self-inverse order automorphism of $\left\langle X,\leqslant_{X}\right\rangle$. This poset is drawn in Figure 1.
[…]
Figure 1: The poset $\langle X,\leqslant_{X}\rangle$ from a chain $\langle C,\leqslant\rangle$ with order automorphism $\alpha:X\to X$.
[…]
Lemma 3.1
Let $\langle C,\leqslant\rangle$ be a chain, and $\langle X,\leqslant_{X}\rangle$ with $\alpha$ as above. Then $\leqslant_{X}\,=0=(\leqslant_{X})^{c\smile}\circ\alpha$.
Proof 3.2
To prove the left-to-right inclusion, assume $x^{b}\leqslant_{X}y^{d}$. Then $x^{b}=y^{d}$ or $x<y$. If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$. Hence, $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, i.e., $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}\circ\alpha=0$. On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$. Again we get $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}=0$.
For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$. Then we have $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, which means $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$. We therefore get $y^{-d}\not\leqslant x^{b}$. If $b=-d$, then it must be the case that $y\not\leqslant x$, in which case we have $x<y$. Hence, $x^{b}\leqslant_{X}y^{d}$. On the other hand, if $b\neq-d$, then $y^{-d}\not\leqslant x^{b}$ implies $y\not<x$. This means $x\leqslant y$, and so, since $d=b$, we get $x^{b}\leqslant_{X}y^{d}$.
Consider the rationals $\left\langle\mathbb{Q},\leqslant\right\rangle$ with their usual order. Let $\mathbb{Q}^{n}=\left\{\left(q_{1},\ldots,q_{n}\right)\mid q_{1},\ldots,q_{n}\in\mathbb{Q}\right\}$. We will again denote the $n$-tuple $\left(q_{1},\dots,q_{n}\right)$ by $q$. Define binary relations $\mathrm{id}_{\mathbb{Q}^{n}}$ and $L_{1}$ on $\mathbb{Q}^{n}$ by
[…]
and for $i\in\left\{2,\ldots,n\right\}$, define binary relations $L_{i}$ on $\mathbb{Q}^{n}$ by
[…]
Now set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$.
The lemma below will help to simplify some of the proofs later on.
Lemma 3.3
Let $n\geqslant 1$ and $j,k\in\left\{1,\ldots,n\right\}$. If $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$, then $\left(p,r\right)\in L_{\textnormal{min}\{j,k\}}$.
Proof 3.4
Let $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$. Then $\left(p_{1},\ldots,p_{j-1}\right)=\left(q_{1},\ldots,q_{j-1}\right)$, $p_{j}<_{\mathbb{Q}}q_{j}$, $\left(q_{1},\ldots,q_{k-1}\right)=\left(r_{1},\ldots,r_{k-1}\right)$ and $q_{k}<_{\mathbb{Q}}r_{k}$. Clearly then $\left(p_{1},\ldots,p_{\textnormal{min}\{j,k\}-1}\right)=\left(r_{1},\ldots,r_{\textnormal{min}\{j,k\}-1}\right)$. If $j=k$, then we get $p_{j}<q_{j}<r_{j}$. If $\text{min}\{j,k\}=j$, then $p_{j}<q_{j}=r_{j}$. Finally, if $\text{min}\{j,k\}=k$, then $p_{k}=p_{k}=q_{k}<r_{k}$. Hence, in all three cases we have $(p,r)\in L_{\text{min}\{j,k\}}$.
Lemma 3.5
Let $n\geqslant 1$. Then $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
Proof 3.6
The relation $\leqslant_{n}$ is clearly reflexive and anti-symmetric. For transitivity, assume $p\leqslant_{n}q$ and $q\leqslant_{n}r$. If $p=q$ or $q=r$, then we clearly have $p\leqslant_{n}r$. Assume there are $j,k\in\left\{1,\ldots,n\right\}$ such that $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$. In this case we have $\left(p,r\right)\in L_{{\textnormal{min}\{j,k\}}}$ by Lemma 3.3, which means $p\leqslant_{n}r$.
Finally, we show that $\leqslant_{n}$ is a linear order. If $p=q$, we are done, so suppose $p\neq q$. Then there is a smallest $i\in\left\{1,\ldots,n\right\}$ such that $p_{i}\neq q_{i}$. Since the set of rational numbers is linearly ordered, we have $p_{i}<_{\mathbb{Q}}q_{i}$ or $q_{i}<_{\mathbb{Q}}p_{i}$. If $i=1$, then $\left(p,q\right)\in L_{1}$ or $\left(q,p\right)\in L_{1}$, which means $p\leqslant_{n}q$ or $q\leqslant_{n}p$. On the other hand, if $i\in\{2,\ldots,n\}$, then $p_{j}=q_{j}$ for all $j\in\left\{1,\ldots,i-1\right\}$. Hence, $\left(p,q\right)\in L_{i}$ or $\left(q,p\right)\in L_{i}$, and therefore $p\leqslant_{n}q$ or $q\leqslant_{n}p$.
Let $X:=\left(\mathbb{Q}^{n}\times\left\{-1\right\}\right)\cup\left(\mathbb{Q}^{n}\times\left\{1\right\}\right)$, and define a binary relation $\leqslant_{X}$ on $X$ by
$p^{b}\leqslant_{X}q^{d}\quad\text{ iff }\quad p^{b}=q^{d}\quad\textnormal{ or }\quad p<_{n}q$
Consider the posets $\left\langle X,\leqslant_{X}\right\rangle$ and $\left\langle X^{2},\preccurlyeq\right\rangle$ (see Section 2.3). Let $\alpha:X\to X$ be defined by $\alpha\left(p^{b}\right)=p^{-b}$. For each $1\leqslant j\leqslant n$, set
$U_{j}=\left\{\left(p^{b},q^{d}\right)\Big{|}\;b,d\in\left\{-1,1\right\}\text{ and }\left(p,q\right)\in L_{j}\right\}.$
Using $\alpha$ and $U_{j}$, we alter the definition of the binary relation $T_{i}$ in [16, Theorem 6.2] and Example 2.11 as follows. For $i\in\left\{-n-1,\ldots,n+1\right\}$, define
$R_{i}=\begin{cases}\varnothing&\text{if }\,i=-n-1\\ \displaystyle\bigcup_{j=1}^{n+1+i}U_{j}&\text{if }\,-n\leqslant i\leqslant-1\\ \leqslant_{X}&\text{if }\,i=0\\ {\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha&\text{if }\,1\leqslant i\leqslant n+1\end{cases}$
The relations $R_{i}$ will be the elements of our representation of the odd Sugihara monoid $\mathbf{S}_{2n+3}$.
To see how the construction of Maddux [16] (see Example 2.11) relates to our construction, consider the map $\delta$ that sends relations $R$ on $\left(\mathbb{Q}^{n}\right)^{2}$ to relations on $X$ as follows:
$\delta\left(R\right)=\left\{\left(p^{b},q^{d}\right)\mid b,d\in\{-1,1\}\textnormal{ and }\left(p,q\right)\in R\right\}.$
Clearly $\delta\left(L_{j}\right)=U_{j}$. It is straightforward to show that
$\delta\left(\bigcup_{j=1}^{n+1+i}L_{j}\right)=\bigcup_{j=1}^{n+1+i}\delta\left(L_{j}\right)$
which means $\delta\left(T_{i}\right)=R_{i}$ for all $i\in\left\{-n-1,\ldots,-1\right\}$. Moreover, we can also show that $\delta\left(\left(T_{-i}\right)^{c\smile}\right)=\left(\delta\left(T_{-i}\right)\right)^{c\smile}$ for all $i\in\left\{1,\ldots,n+1\right\}$, which implies that $\delta\left(T_{i}\right)\circ\alpha=R_{i}$ for all $i\in\left\{1,\ldots,n+1\right\}$.
Lemma 3.7
For each $i\in\left\{-n-1,\ldots,n+1\right\}$, the relation $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$.
Proof 3.8
Both $X^{2}$ and $\varnothing$ are upsets of $\left\langle X^{2},\preccurlyeq\right\rangle$. The fact that $\leqslant_{X}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ follows from Lemma 2.4. It would also follow from Lemmas 2.3 and 2.6 that $R_{i}$ is an upset for $i\in\left\{1,\ldots,n\right\}$, if we can show that $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ when $-n\leqslant i<0$.
Let $(p^{b},q^{d})\in R_{i}$ and assume $\left(p^{b},q^{d}\right)\preccurlyeq\left(r^{e},s^{f}\right)$. The former implies that $(p^{e},q^{f})\in U_{j}$ for some $j$ such that $1\leqslant j\leqslant n+1+i$. Hence, $(p,q)\in L_{j}$ for some $j$ such that $1\leqslant j\leqslant n+1+i$. If $j=1$, then we have $p_{1}<q_{1}$. But $\left(p^{b},q^{d}\right)\preccurlyeq\left(r^{e},s^{f}\right)$ implies that $r\leqslant_{n}p$ and $q\leqslant_{n}s$, so $r_{1}\leqslant p_{1}<q_{1}\leqslant s_{1}$. Hence, $\left(r^{e},s^{f}\right)\in U_{1}\subseteq R_{i}$.
On the other hand, if $1<j\leqslant n+1+i$, then $\left(p_{1},\ldots,p_{j-1}\right)=\left(q_{1},\ldots,q_{j-1}\right)$ and $p_{j}<q_{j}$. We now consider the following cases:
Case 1: $r=p$ and $q=s$. In this case we get $\left(r_{1},\ldots,r_{j-1}\right)=\left(p_{1},\ldots,p_{j-1}\right)=\left(q_{1},\ldots,q_{j-1}\right)=\left(s_{1},\ldots,s_{j-1}\right)$ and $r_{j}=p_{j}<q_{j}=s_{j}$. Hence, $\left(r^{e},s^{f}\right)\in U_{j}\subseteq R_{i}$.
Case 2: $r_{1}<p_{1}$ or $q_{1}<s_{1}$. Then $r_{1}<p_{1}=q_{1}\leqslant s_{1}$ or $r_{1}\leqslant p_{1}=q_{1}<s_{1}$. In both cases, we have $\left(r^{e},s^{f}\right)\in U_{1}\subseteq R_{i}$.
Case 3: $r=p$ and $\left(q,s\right)\in L_{k}$ for some $k\in\left\{2,\ldots,n\right\}$. Here we have $\left(p,s\right)\in L_{\textnormal{min}\{j,k\}}$ by Lemma 3.3. Hence, since $r=p$, it follows that $\left(r_{1},\ldots,r_{\textnormal{min}\{j,k\}-1}\right)=\left(p_{1},\ldots,p_{\textnormal{min}\{j,k\}-1}\right)$ and $r_{\textnormal{min}\{j,k\}}=p_{\textnormal{min}\{j,k\}}<s_{\textnormal{min}\{j,k\}}$. This means $\left(r,s\right)\in L_{\textnormal{min}\{j,k\}}$, and so $\left(r^{e},s^{f}\right)\in U_{\text{min}\{j,k\}}\subseteq R_{i}$.
Case 4: $q=s$ and $\left(r,p\right)\in L_{k}$ for some $k\in\left\{2,\ldots,n\right\}$. This case is similar to Case 3.
Case 5: $\left(r,p\right)\in L_{k}$ and $\left(q,s\right)\in L_{\ell}$ for $k,\ell\in\left\{2,\ldots,n\right\}$. By Lemma 3.3, $\left(r,q\right)\in L_{\text{min}\{j,k\}}$, and so, by another application of Lemma 3.3, we have $\left(r,s\right)\in L_{\text{min}\{\text{min}\{j,k\},\ell\}}$. Hence, $\left(r,s\right)\in L_{\text{min}\{j,k,\ell\}}$, which means $\left(r^{e},s^{f}\right)\in U_{\text{min}\{j,k,\ell\}}\subseteq R_{i}$.
Notice that when defining $R_{i}$ for $i\in\left\{1,\ldots,n+1\right\}$ we chose to use the unary order-reversing operation $\sim$ instead of --. Since $\alpha\neq\text{id}_{X}$ we know from Theorem 2.7 that the algebra will not be cyclic and hence ${\sim}R\neq-R$ in general. However, the following lemma shows that for the relations $R_{i}$, we will in fact get ${\sim}R_{i}=-R_{i}$.
Lemma 3.9
For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $-R_{i}={\sim}R_{i}$.
Proof 3.10
If $i=0$, the claim follows from Lemma 2.5. Now consider the following cases:
Case 1: $-n-1\leqslant i<0$. Let $\left(p^{b},q^{d}\right)\notin-R_{i}$. By Lemma 2.2, $\left(p^{b},q^{d}\right)\in\left(-R_{i}\right)^{c}=\left(\alpha\circ\left(R_{i}\right)^{c\smile}\right)^{c}=\alpha\;\circ\;R^{\smile}_{i}$. Therefore, $\left(q^{d},p^{-b}\right)\in R_{i}$, and so $\left(q^{d},p^{-b}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. This implies $(q,p)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. It thus follows that $\left(q^{-d},p^{b}\right)\in U_{k}\subseteq R_{i}$, i.e., $\left(\alpha\left(q^{d}\right),p^{b}\right)\in U_{k}\subseteq R_{i}$. Hence, $\left(p^{b},\alpha\left(q^{d}\right)\right)\in R_{i}^{\smile}$, and therefore $\left(p^{b},\alpha\left(q^{d}\right)\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}$. From this we get $\left(p^{b},q^{d}\right)\in\left(\left(R_{i}\right)^{c\smile}\right)^{c}\circ\alpha=\left(\left(R_{i}\right)^{c\smile}\circ\alpha\right)^{c}=\left({\sim}R_{i}\right)^{c}$, i.e., $\left(p^{b},q^{d}\right)\notin{\sim}R_{i}$.
The reverse inclusion can be proved in a similar way.
Case 2: $10<i\leqslant n+$ 1. This case follows from the previous case and the fact that -- and $\sim$ satisfy the involutive property.
We want our set of relations to have the structure of a Sugihara chain, therefore we need to confirm that they form a linear order when ordered by inclusion.
Lemma 3.11
The relations in $\left\{R_{-n-1},R_{-n},\ldots,R_{-1},R_{0},R_{1},\ldots,R_{n},R_{n+1}\right\}$ form a chain, i.e.,
$R_{-n-1}\subseteq R_{-n}\subseteq\cdots\subseteq R_{-1}\subseteq R_{0}\subseteq R_{1}\subseteq\cdots\subseteq R_{n}\subseteq R_{n+1}.$
Proof 3.12
It follows from the construction that $R_{-i}\subseteq R_{-i+1}$ for all $i\in\left\{2,\ldots,n+1\right\}$. Since $\sim$ is order-reversing, it follows that $R_{i}\subseteq R_{i+1}$ for all $i\in\left\{1,\ldots,n\right\}$. To see that $R_{-1}\subseteq R_{0}=\;\leqslant_{X}$, let $\left(p^{b},q^{d}\right)\in R_{-1}$. Then $\left(p^{b},q^{d}\right)\in U_{j}$ for some $j\in\left\{1,\ldots,n\right\}$, and therefore $(p,q)\in L_{j}$ for some $j\in\left\{1,\ldots,n\right\}$. Hence, $p<_{n}q$, which means $p^{b}\leqslant_{X}q^{d}$. The inclusion $R_{0}\subseteq R_{1}$ then follows from the fact that $\sim$ is order-reversing.
In $\delta$-notation, the above result can be rewritten as follows:
$\delta\left(T_{-n-1}\right)\subseteq\delta\left(T_{-n}\right)\subseteq\cdots\subseteq\delta\left(T_{-1}\right)\subseteq R_{0}\subseteq\delta\left(T_{1}\right)\circ\alpha\subseteq\cdots\subseteq\delta\left(T_{n}\right)\circ\alpha\subseteq\delta\left(T_{n+1}\right)\circ\alpha$
The transitivity of each of the relations $R_{i}$ will be essential for the calculations involving relational composition.
Lemma 3.13
For each $i\in\left\{-n-1,\ldots,n+1\right\}$, the relation $R_{i}$ is transitive.
Proof 3.14
This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for $i=-n-1$. We now consider the following two cases:
Case 1: $-n\leqslant i<0$. Let $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. Then there exist $j,k\in\left\{1,\ldots,n+1+i\right\}$ such that $(p,r)\in L_{j}$ and $(r,q)\in L_{k}$. Hence, by Lemma 3.3, we have $\left(p,q\right)\in L_{\text{min}\{j,k\}}$, and so $\left(p^{b},q^{d}\right)\in U_{\text{min}\{j,k\}}\subseteq R_{i}$.
Case 2: $0<i\leqslant n$. Let $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. Then we have $\left(p^{b},r^{e}\right)\in{\sim}R_{-i}$ and $\left(r^{e},q^{d}\right)\in{\sim}R_{-i}$. Hence, $\left(r^{-e},p^{b}\right)\notin R_{-i}$ and $\left(q^{-d},r^{e}\right)\notin R_{-i}$, which means $(r,p)\notin L_{j}$ for all $j\in\left\{1,\ldots,n+1-i\right\}$ and $(q,r)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. The first part implies that $r_{1}\not<p_{1}$ and, for all $j\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{j-1}\right)=\left(p_{1},\ldots,p_{j-1}\right)$ then $r_{j}\not<p_{j}$, while the second part gives $q_{1}\not<r_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(q_{1},\ldots,q_{k-1}\right)=\left(r_{1},\ldots,r_{k-1}\right)$ then $q_{k}\not<r_{k}$. Since $\mathbb{Q}$ is linearly ordered, $p_{1}\leqslant r_{1}$ and, for all $j\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{j-1}\right)=\left(p_{1},\ldots,p_{j-1}\right)$ then $p_{j}\leqslant r_{j}$. Likewise, $r_{1}\leqslant q_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(q_{1},\ldots,q_{k-1}\right)=\left(r_{1},\ldots,r_{k-1}\right)$ then $r_{k}\leqslant q_{k}$. If $p_{1}<r_{1}$, then we have $p_{1}<r_{1}\leqslant q_{1}$, and so $(p,q)\in L_{1}$, which means $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{-1}\subseteq R_{i}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, so again $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{-1}\subseteq R_{i}$. Finally, suppose $p_{1}=r_{1}=q_{1}$. There are four cases:
  1. (a)
    There exist $\ell,m\in\left\{2,\ldots,n+1-i\right\}$ such that $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$, $p_{\ell}<r_{\ell}$, $\left(q_{1},\ldots,q_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)$ and $r_{m}<q_{m}$; that is, $\left(p,r\right)\in L_{\ell}$ and $\left(r,q\right)\in L_{m}$. Therefore, by Lemma 3.3, we get $\left(p,q\right)\in L_{\textnormal{min}\{\ell,m\}}$, and thus $\left(p^{b},q^{d}\right)\in U_{\textnormal{min}\{\ell,m\}}\subseteq R_{-i}\subseteq R_{i}$.
  2. (b)
    There exists $\ell\in\left\{2,\ldots,n+1-i\right\}$ such that $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ and $p_{\ell}<r_{\ell}$, and for all $k\in\left\{2,\ldots,n+1-i\right\}$, we have $r_{k}=q_{k}$. In this case, $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ and $p_{\ell}<r_{\ell}=q_{\ell}$. Hence, $(p,q)\in L_{\ell}$, which gives $\left(p^{b},q^{d}\right)\in U_{\ell}\subseteq R_{-i}\subseteq R_{i}$.
  3. (c)
    There exists $m\in\left\{2,\ldots,n+1-i\right\}$ such that $\left(q_{1},\ldots,q_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)$ and $r_{m}<q_{m}$, and for all $j\in\left\{1,\ldots,n+1-i\right\}$, we have $p_{j}=r_{j}$. Then $\left(p_{1},\ldots,p_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}=r_{m}<q_{m}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{-i}\subseteq R_{i}$.
  4. (d)
    For all $j,k\in\left\{1,\ldots,n+1-i\right\}$, we have $p_{j}=r_{j}$ and $q_{k}=r_{k}$. This implies $(q,p)\notin L_{j}$ for all $j\in\left\{1,\ldots,n+1-i\right\}$, so $\left(q^{-d},p^{b}\right)\notin U_{j}$ for all $j\in\left\{1,\ldots,n+1-i\right\}$. Hence, it follows that $\left(q^{-d},p^{b}\right)\notin R_{-i}$, which means $\left(p^{b},q^{-d}\right)\in\left(R_{-i}\right)^{c\smile}$, i.e., $\left(p^{b},\alpha\left(q^{d}\right)\right)\in\left(R_{-i}\right)^{c\smile}$. We thus get $\left(p^{b},q^{d}\right)\in\left(R_{-i}\right)^{c\smile}\circ\alpha={\sim}R_{-i}=R_{i}$.
Given $n\geqslant 1$, in order for the relations $\{\,R_{i}\mid-(n+1)\leqslant i\leqslant n+1\,\}$ to give a representation of the Sugihara chain $\mathbf{S}_{2n+3}$, we need to show the following:
$R_{i}\circ R_{j}=\begin{cases}R_{i}&\text{if }\,|j|<|i|\\ R_{j}&\text{if }\,|i|<|j|\\ R_{\text{min}\{i,j\}}&\text{if }\,|j|=|i|\end{cases}$
We will use a series of lemmas to show that the above holds. Each lemma will involve a number of cases. We will frequently make use of the fact that composition is order-preserving in both coordinates, and that in all cases the relations are greater than or less than the monoid identity $\leqslant_{X}$.
Lemma 3.15
If $|i|>|j|$ then $R_{i}\circ R_{j}=R_{i}$.
Proof 3.16
If $i=-n-1$ (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
Case 1: $-n-1<i<j<0$. Since $R_{j}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Hence, there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p^{b},q^{d}\right)\in U_{k}$, i.e., there exists $k\in\left\{1,\ldots,n+1+i\right\}$ such that $\left(p,q\right)\in L_{k}$. There are two cases:
  1. (a)
    $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there is some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Let $r=(r_{1},0,\ldots,0)$. Then $\left(p,r\right)\in L_{1}$ and $(r,q)\in L_{1}$, i.e., $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{i}\subseteq R_{j}$. We thus have $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
  2. (b)
    There is some $k\in\left\{2,\ldots,n+1+i\right\}$ such that $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$. By the density of $\mathbb{Q}$, there exists $r_{k}$ such that $p_{k}<r_{k}<q_{k}$. Let $r=(p_{1},\dots,p_{k-1},r_{k},0,\dots,0)$. Clearly $(p^{b},r^{d})\in U_{k}\subseteq R_{i}$. Now notice that $r=(q_{1},\dots,q_{k-1},r_{k},0,\dots,0)$, and therefore $(r^{d},q^{d})\in U_{k}\subseteq R_{i}\subseteq R_{j}$. This gives us $(p^{b},q^{d})\in R_{i}\circ R_{j}$ and so $R_{i}\circ R_{j}=R_{i}$.
Case 2: $-n-1<i<0$, $j>0$. Since ${\leqslant_{X}}\subseteq R_{j}$, we get $R_{i}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. From the first part we get $\left(p^{b},r^{e}\right)\in U_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$, which means $(p,r)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The second part gives $\left(r^{e},q^{d}\right)\in{\sim}R_{-j}=\left(R_{-j}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},r^{b}\right)\notin R_{-j}$. This implies that $\left(q^{-d},r^{e}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. That is, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{1}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, and so $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $r_{1}<q_{1}$, we have $p_{1}\leqslant r_{1}<q_{1}$, so again $\left(p,q\right)\in L_{1}$, i.e., $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}$. If $p_{1}=r_{1}=q_{1}$, then it must be the case that $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}$.
Case 3: $i>0$, $-n-1<j<0$. Since $R_{j}\subseteq{\leqslant_{X}}$ we get $R_{i}\circ R_{j}\subseteq R_{i}$. Let $\left(p^{b},q^{d}\right)\in R_{i}$. Then $\left(p^{b},q^{d}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},p^{b}\right)\notin R_{-i}$. This means that $\left(q^{-d},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. That is, $q\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $q_{k}\not<p_{k}$. We therefore obtain $p_{1}\leqslant q_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{k-1}\right)=\left(q_{1},\ldots,q_{k-1}\right)$ then $p_{k}\leqslant q_{k}$. Suppose $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there is some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Now let $r=\left(p_{1},0,\ldots,0\right)$. Then $(p,r)\in L_{1}$ and $(r,q)\in L_{1}$, i.e., $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{-i}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. On the other hand, suppose $p_{1}=q_{1}$. We have two cases:
  1. (a)
    There is some $\ell\in\left\{2,\ldots,n+1-i\right\}$ such that $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ and $p_{\ell}<q_{\ell}$. Then $\left(p,q\right)\in L_{\ell}$, and so $\left(p^{b},q^{d}\right)\in U_{\ell}\subseteq R_{-i}\subseteq R_{j}$. Since $\leqslant_{X}\,\subseteq R_{i}$, we have $\left(p^{b},p^{b}\right)\in R_{i}$, and therefore we obtain $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
  2. (b)
    For all $k\in\left\{2,\ldots,n+1-i\right\}$, we have $p_{k}=q_{k}$. Let $r=\left(q_{1},\ldots,q_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$. Then $(r,q)\in L_{n+2-i}$, and so $\left(r^{b},q^{d}\right)\in U_{n+2-i}\subseteq\bigcup_{k=1}^{n+1+j}\subseteq R_{j}$. Note now that we also have $r=\left(p_{1},\ldots,p_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$, so $\left(r^{-b},p^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $\left(r^{-b},p^{b}\right)\notin R_{-i}$, which implies that $\left(p^{b},r^{-b}\right)\in\left(R_{-i}\right)^{c\smile}$. It therefore follows that $\left(p^{b},r^{b}\right)\in\left(R_{-i}\right)^{c\smile}\circ\alpha={\sim}R_{-i}=R_{i}$. From this we get $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Case 4: $i>j>0$. Since ${\leqslant_{X}}\subseteq R_{j}$ we get $R_{i}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. But $R_{j}\subseteq R_{i}$ and $R_{i}$ is transitive by Lemma 3.13, so $\left(p^{b},q^{d}\right)\in R_{i}$.
Lemma 3.17
If $|j|>|i|$ then $R_{i}\circ R_{j}=R_{j}$.
Proof 3.18
If $j=-n-1$, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
Case 1: $-n-1<j<i<0$. Since $R_{i}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{j}$. Then $\left(p,q\right)\in L_{k}$ for some $k\in\left\{1,2,\ldots,n+1+j\right\}$. That is, $p_{1}<q_{1}$ or $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$. If $p_{1}<q_{1}$, we use the density of $\mathbb{Q}$ to find some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Then we set $r=\left(r_{1},0\ldots,0\right)$ and obtain $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Similarly, if $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$ for some $k\in\left\{2,\ldots,n+1+j\right\}$, we can find some $r_{k}\in\mathbb{Q}$ such that $p_{k}<r_{k}<q_{k}$. Setting $r=\left(q_{1},\ldots,q_{k-1},r_{k},0,\ldots,0\right)$, we obtain $\left(p^{b},r^{b}\right)\in U_{k}\subseteq R_{j}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{k}\subseteq R_{j}$, which means $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Case 2: $i>0,-n-1<j<0$. Since ${\leqslant_{X}}\subseteq R_{i}$, we get $R_{j}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The latter gives $(r,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+j\right\}$. The former gives $\left(p^{b},r^{e}\right)\in{\sim}R_{-i}=\left(R_{-i}\right)^{c\smile}\circ\alpha$, and so $\left(r^{-e},p^{b}\right)\notin R_{-i}$. Hence, $\left(r^{-e},p^{b}\right)\notin U_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-i\right\}$. That is, $r_{1}\not<p_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-i\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $r_{\ell}\not<p_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(r_{1},\ldots,r_{\ell-1}\right)$ then $p_{\ell}\leqslant r_{1}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}<r_{1}$, we get $p_{1}<r_{1}\leqslant q_{1}$, so again $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. If $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}$.
Case 3: $i<0,j>0$. Since $R_{i}\subseteq{\leqslant_{X}}$, we have $R_{i}\circ R_{j}\subseteq R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{j}$. Then $\left(p^{b},q^{d}\right)\in{\sim}R_{-j}=\left(R_{-j}\right)^{c\smile}\circ\alpha$, and so $\left(q^{-d},p^{b}\right)\notin R_{-j}$. This means that $\left(q,p\right)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$. Hence, $q_{1}\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $q_{k}\not<p_{k}$. Since $\mathbb{Q}$ is linearly ordered, $p_{1}\leqslant q_{1}$ and, for all $k\in\left\{2,\ldots,n+1-j\right\}$, if $\left(q_{1},\ldots,q_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $p_{k}\leqslant q_{k}$. Suppose $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there is some $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Let $r=\left(r_{1},0,\ldots,0\right)$. Then we have $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{-j}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{-j}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Now suppose $p_{1}=q_{1}$. Then there are two cases:
  1. (a)
    There is some $\ell\in\left\{2,\ldots,n+1-j\right\}$ such that $\left(p_{1},\ldots,p_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ and $p_{\ell}<q_{\ell}$. Then we have $\left(p,q\right)\in L_{\ell}$, and so $\left(p^{b},q^{d}\right)\in U_{\ell}\subseteq R_{-j}\subseteq R_{i}$. But $\left(q^{d},q^{d}\right)\in R_{j}$ since $\leqslant_{X}\,\subseteq R_{j}$, so it follows that $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
  2. (b)
    For all $k\in\left\{1,\ldots,n+1-j\right\}$, we have $p_{k}=q_{k}$. Let
    $r=\left(p_{1},\ldots,p_{n+1-j},p_{n+2-j}+1,\ldots,p_{n}+1\right).$
    Then $\left(p,r\right)\in L_{n+2-j}$, and so $\left(p^{b},r^{b}\right)\in U_{n+2-j}\subseteq\bigcup_{k=1}^{n+1+i}U_{k}=R_{i}$. We also have $\left(q,r\right)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$. Hence, $\left(q^{-d},r^{b}\right)\notin U_{k}$ for all $k\in\left\{1,\ldots,n+1-j\right\}$, which means $\left(q^{-d},p^{b}\right)\notin R_{-j}$. We thus have $\left(r^{b},q^{-d}\right)\in\left(R_{-j}\right)^{c\smile}$, and thus, $\left(r^{b},q^{d}\right)\in\left(R_{-j}\right)^{c\smile}\circ\alpha={\sim}R_{-j}=R_{j}$. This gives us $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Case 4: $j>i>0$. Since ${\leqslant_{X}}\subseteq R_{i}$ we have $R_{j}\subseteq R_{i}\circ R_{j}$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$. Then there is some $r^{e}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. But $R_{i}\subseteq R_{j}$ and $R_{j}$ is transitive by Lemma 3.13, so $\left(p^{b},q^{d}\right)\in R_{j}$.
Lemma 3.19
If $|j|=|i|$ then $R_{i}\circ R_{j}=R_{\text{min}\{i,j\}}$.
Proof 3.20
The claim clearly follows for the cases $i=j=0$ and $i=j=-n-1$. So consider the following cases:
Case 1: $-n-1<i=j<0$. Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}=R_{i}\circ R_{i}$. Then there exists $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$. By Lemma 3.13, $R_{i}$ is transitive, so $\left(p^{b},q^{d}\right)\in R_{i}=R_{j}=R_{\text{min}\{i,j\}}$. Now let $\left(p^{b},q^{d}\right)\in R_{\text{min}\{i,j\}}=R_{i}=R_{j}$. Then $(p,q)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. There are two cases:
  1. (a)
    $p_{1}<q_{1}$. By the density of $\mathbb{Q}$, there exists $r_{1}\in\mathbb{Q}$ such that $p_{1}<r_{1}<q_{1}$. Let $r=\left(r_{1},0,\ldots,0\right)$. Then $\left(p^{b},r^{b}\right)\in U_{1}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{1}\subseteq R_{j}$. Hence, $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
  2. (b)
    There exists $k\in\left\{2,\ldots,n+1+i\right\}$ such that $(p_{1},\dots,p_{k-1})=(q_{1},\dots,q_{k-1})$ and $p_{k}<q_{k}$. By the density of $\mathbb{Q}$, there exists $r_{k}$ such that $p_{k}<r_{k}<q_{k}$. Set $r=(p_{1},\dots,p_{k-1},r_{k},0,\dots,0)$. Then $(p^{b},r^{d})\in U_{k}\subseteq R_{i}$ and $(r^{d},q^{d})\in U_{k}\subseteq R_{j}$. This gives us $(p^{b},q^{d})\in R_{i}\circ R_{j}$.
Case 2: $-n-1<i<0$, $j>0$. Since $\leqslant_{X}\;\subseteq R_{j}$, we have $R_{\text{min}\{i,j\}}=R_{i}\subseteq R_{i}\circ R_{j}$. Now let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there is some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The former gives $\left(p,r\right)\in L_{k}$ for some $k\in\left\{1,\ldots,n+1+i\right\}$. The latter gives $\left(q^{-d},r^{e}\right)\notin R_{-j}$, and so $(q,r)\notin L_{\ell}$ for all $\ell\in\left\{1,\ldots,n+1-j\right\}$. Hence, $q_{1}\not<r_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $q_{\ell}\not<r_{\ell}$. Since $\mathbb{Q}$ is linearly ordered, we obtain $r_{1}\leqslant q_{1}$ and, for all $\ell\in\left\{2,\ldots,n+1-j\right\}$, if $\left(r_{1},\ldots,r_{\ell-1}\right)=\left(q_{1},\ldots,q_{\ell-1}\right)$ then $r_{\ell}\leqslant q_{\ell}$. If $p_{1}<r_{1}$, then $p_{1}<r_{1}\leqslant q_{1}$, so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. If $r_{1}<q_{1}$, we get $p_{1}\leqslant r_{1}<q_{1}$, which implies $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{i}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,k\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}=R_{\text{min}\{i,j\}}$.
Case 3: $-n-1<j<0,i>0$. Since $\leqslant_{X}\;\subseteq R_{i}$, we have $R_{\text{min}\{i,j\}}=R_{j}\subseteq R_{i}\circ R_{j}$. Let $(p^{b},q^{d})\in R_{i}\circ R_{j}$. Then there exists some $r^{b}$ such that $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{j}$. The second part gives $\left(r,q\right)\in L_{\ell}$ for some $\ell\in\left\{1,\ldots,n+1+j\right\}$. The first part gives $\left(r^{-e},p^{b}\right)\notin R_{-i}$, which means $(r,p)\notin L_{k}$ for all $k\in\left\{1,\ldots,n+1-i\right\}$. Hence, $r_{1}\not<p_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $r_{k}\not<p_{k}$. Since $\mathbb{Q}$ is linearly ordered, we get $p_{1}\leqslant r_{1}$ and, for all $k\in\left\{2,\ldots,n+1-i\right\}$, if $\left(r_{1},\ldots,r_{k-1}\right)=\left(p_{1},\ldots,p_{k-1}\right)$ then $p_{k}\leqslant r_{k}$. If $r_{1}<q_{1}$, then $p_{1}\leqslant r_{1}<q_{1}$, and so $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. If $p_{1}<r_{1}$, then it follows that $p_{1}<r_{1}\leqslant q_{1}$, which gives $\left(p^{b},q^{d}\right)\in U_{1}\subseteq R_{j}=R_{\text{min}\{i,j\}}$. Finally, if $p_{1}=r_{1}=q_{1}$, then $\left(p_{1},\ldots,p_{m-1}\right)=\left(r_{1},\ldots,r_{m-1}\right)=\left(q_{1},\ldots,q_{m-1}\right)$ and $p_{m}<q_{m}$ for some $m\in\left\{2,\ldots,\ell\right\}$. Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{j}=R_{\text{min}\{i,j\}}$.
Case 4: $i=j>0$. The left-to-right inclusion follows from the transitivity of $R_{i}=R_{j}$. Now let $\left(p^{b},q^{d}\right)\in R_{\text{min}\left\{i,j\right\}}=R_{i}=R_{j}$. Then $\left(p^{b},p^{b}\right)\in\;\leqslant_{X}\,\subseteq R_{i}$, and hence $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
Combining the three lemmas above gives us the following result.
Proposition 3.21
Let $n\geqslant 1$ and $i,j\in\{-(n+1),\dots,n+1\}$. Then
$R_{i}\circ R_{j}=\begin{cases}R_{i}&\text{if }\,|j|<|i|\\ R_{j}&\text{if }\,|i|<|j|\\ R_{\text{min}\{i,j\}}&\text{if }\,|j|=|i|\end{cases}$
Let $n\geqslant 1$ and consider $\mathcal{R}_{2n+3}=\{\,R_{i}\mid-(n+1)\leqslant i\leqslant n+1\,\}$. Then define the algebra
$\mathbb{S}_{2n+3}:=\langle\mathcal{R}_{2n+3},\cap,\cup,\circ,\Rightarrow,R_{0},{\sim}\rangle$
[…]
Theorem 3.22
Let $n\geqslant 1$. Then the algebra $\mathbb{S}_{2n+3}$ is an odd Sugihara chain and $\mathbf{S}_{2n+3}\cong\mathbb{S}_{2n+3}$.
Proof 3.23
[…]
4 Representing all Sugihara monoids via weakening relations
The main result of this section is to show that all Sugihara monoids can be represented as algebras of binary relations with the monoid operation given by relational composition. Since the algebras of relations are constructed using the method of Section 2.3, the relations will be weakening relations. We begin with an important definition.
Definition 4.1
A Sugihara monoid $\mathbf{A}$ is representable if it is isomorphic to the direct reduct of an element of $\mathsf{RDInFL}$.
The class of representable Sugihara monoids will be denoted by $\mathsf{RSM}$. We note that since $\mathsf{RDInFL}$ is closed under $\mathbb{I}$, $\mathbb{S}$ and $\mathbb{P}$ (see Definition 2.14), we will also have that $\mathsf{RSM}=\mathbb{ISP}(\mathsf{RSM})$.
From here, we proceed as follows. We show that the class of representable DInFL-algebras is closed under ultraproducts. Then, using the closure of $\mathsf{RDInFL}$ under ultraproducts, we are able to show that the Sugihara monoids $\mathbf{S}$ and $\mathbf{S}^{*}$ are representable. Since those two algebras generate the quasivariety $\mathsf{SM}$, we will then be able to conclude that $\mathsf{SM}\subseteq\mathsf{RSM}$.
4.1 Closure of $\mathsf{RDInFL}$ under ultraproducts
Here we show that the class of representable DInFL-algebras (see Definition 2.14) is closed under taking ultraproducts. The ultraproduct was first introduced by Łoś [15] in 1955. The construction is now very well-known, both for algebras and other first-order structures (cf. [2, IV-6 and V-2]).
Let $I$ be a set and let $\left\{A_{i}\mid i\in I\right\}$ be a family of non-empty sets indexed by $I$. Recall that if $a,b\in\prod\{A_{i}\mid i\in I\}$, then the subset
$\llbracket a=b\rrbracket=\left\{i\in I\mid a(i)=b(i)\right\}$
of $I$ is called the equaliser of $a$ and $b$.
If $\mathcal{F}$ is a filter on $I$, then we may define a binary relation $\theta_{\mathcal{F}}$ on the Cartesian product $\prod\left\{A_{i}\mid i\in I\right\}$ by setting, for all $a,b\in\prod\left\{A_{i}\mid i\in I\right\}$,
$\left(a,b\right)\in\theta_{\mathcal{F}}\qquad\textnormal{ iff }\qquad\llbracket a=b\rrbracket\in\mathcal{F}.$
It is straightforward to check that $\theta_{\mathcal{F}}$ is an equivalence relation on $\prod\left\{A_{i}\mid i\in I\right\}$.
Lemma 4.2
Let $I$ be a set and let $\mathbf{A}_{i}$ be an algebra, for each $i\in I$. For each filter $\mathcal{F}$ on $I$, the relation $\theta_{\mathcal{F}}$ is a congruence on the direct product $\prod\left\{\mathbf{A}_{i}\mid i\in I\right\}$.
Let $I$ be a set and $\mathbf{A}_{i}$ an algebra, for each $i\in I$. Recall that if the filter $\mathcal{F}$ in Lemma 4.2 is an ultrafilter on $I$, then the algebra $\prod\left\{\mathbf{A}_{i}\mid i\in I\right\}/\theta_{\mathcal{F}}$ is called the ultraproduct of $\left\{\mathbf{A}_{i}\mid i\in I\right\}$. For a class of algebras $\mathsf{K}$, we denote by $\mathbb{P_{U}}(\mathsf{K})$ the class of all ultraproducts of members of $\mathsf{K}$. The following theorem is immediate.
Theorem 4.3
An ultraproduct of a family of DInFL-algebras is again a DInFL-algebra.
Consider an index set $I$ and let $\mathcal{F}$ be an ultrafilter on $I$. For each index $i\in I$, let $\mathbf{X}_{i}=\left\langle X_{i},\leqslant_{i}\right\rangle$ be a poset and $E_{i}$ an equivalence relation on $X_{i}$ such that $X_{i}\neq\varnothing$ and ${\leqslant_{i}}\subseteq E_{i}$. Consider the Cartesian product $X=\prod_{i\in I}X_{i}$ and the equaliser $\llbracket x=y\rrbracket=\left\{i\in I\mid x(i)=y(i)\right\}$ of $x,y\in X$. Define a binary relation on $X$ by setting, for all $x,y\in X$,
$\left(x,y\right)\in E_{\mathcal{F}}\qquad\textnormal{iff}\qquad\llbracket x=y\rrbracket\in\mathcal{F}.$
Then $E_{\mathcal{F}}$ is an equivalence relation on $X$. Now consider the set of equivalence classes $Y=X/E_{\mathcal{F}}$. Define binary relations $\leqslant_{Y}$ and $E_{Y}$ on $Y$ by setting, for all $[x],[y]\in Y$,
$[x]\leqslant_{Y}[y]\qquad$ $\textnormal{iff}\qquad\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F},$
$\left([x],[y]\right)\in E_{Y}\qquad$ $\textnormal{iff}\qquad\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}.$
Suppose further that for each $i\in I$, the map $\alpha_{i}:X_{i}\to X_{i}$ is an order automorphism of $\mathbf{X}_{i}$ such that $\alpha_{i}\subseteq E_{i}$. Define $\alpha_{Y}:Y\to Y$ by setting, for each $[x]\in Y$,
$\alpha_{Y}\left([x]\right)=\left[\alpha_{X}(x)\right]$
where $\left(\alpha_{X}(x)\right)(i)=\alpha_{i}\left(x(i)\right)$ for all $x\in X$ and $i\in I$. Since each of the $\alpha_{i}$ is bijective, we get the following result.
Lemma 4.4
The map $\alpha_{Y}:Y\to Y$ has an inverse. Moreover, $\alpha^{-1}_{Y}\left([x]\right)=\left[\alpha^{-1}_{X}(x)\right]$ where $\left(\alpha^{-1}_{X}(x)\right)(i)=\alpha_{i}^{-1}\left(x(i)\right)$ for all $x\in X$ and $i\in I$.
The above definitions will be used in the results that follow.
Theorem 4.5
Let $I$ be an index set and $\mathcal{F}$ an ultrafilter on $I$. For each index $i\in I$, let $\mathbf{X}_{i}=\left\langle X_{i},\leqslant_{i}\right\rangle$ be a poset and $E_{i}$ an equivalence relation on $X_{i}$ such that $X_{i}\neq\varnothing$ and ${\leqslant_{i}}\subseteq E_{i}$. Suppose further that for each $i\in I$, the map $\alpha_{i}:X_{i}\to X_{i}$ is an order automorphism of $\mathbf{X}_{i}$ such that $\alpha_{i}\subseteq E_{i}$. Then:
  1. (i)
    $\mathbf{Y}=\left\langle Y,\leqslant_{Y}\right\rangle$ is a poset.
  2. (ii)
    $E_{Y}$ is an equivalence relation on $Y$ such that ${\leqslant_{Y}}\subseteq E_{Y}$.
  3. (iii)
    $\alpha_{Y}$ is an order automorphism of $\mathbf{Y}$ such that $\alpha_{Y}\subseteq E_{Y}$.
  4. (iv)
    If $0=\alpha_{Y}\circ\left(\leqslant_{Y}\right)^{c\smile}=\left(\leqslant_{Y}\right)^{c\smile}\circ\alpha_{Y}$, then $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},0,\sim,-\right\rangle$ is a distributive InFL-algebra.
Proof 4.6
We first prove (i). Let $[x]\in Y$. Since $\leqslant_{i}$ is reflexive for all $i\in I$, we get $\left\{i\in I\mid x(i)\leqslant_{i}x(i)\right\}=I$. Now $\mathcal{F}$ is an ultrafilter, so $I\in\mathcal{F}$. Hence, $\left[x\right]\leqslant_{Y}\left[x\right]$.
Next, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[x]$. Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$. Since $\mathcal{F}$ is closed under taking intersections, it follows that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}.$
Using the fact that $\leqslant_{i}$ is anti-symmetric for all $i\in I$, we can show that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}=\left\{i\in I\mid x(i)=y(i)\right\}.$
Hence we get $\left\{i\in I\mid x(i)=y(i)\right\}\in\mathcal{F}$. This shows that $(x,y)\in E_{\mathcal{F}}$, and hence $[x]=[y]$.
For the transitivity, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[z]$. Then $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$, and so $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$. We know that $\leqslant_{i}$ is transitive for all $i\in I$, so it follows that
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}z(i)\right\}\subseteq\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}.$
Hence, since $\mathcal{F}$ is closed under taking supersets, we have $\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}\in\mathcal{F}$, which means $[x]\leqslant_{Y}[z]$.
We next prove item (ii). Firstly, since $E_{i}$ is reflexive for all $i\in I$ and $I\in\mathcal{F}$, it follows that $\left([x],[x]\right)\in E_{Y}$ for all $[x]\in Y$. Using a similar argument as for item (i), we can show that $E_{Y}$ is transitive. For symmetry, assume $\left([x],[y]\right)\in E_{Y}$. Then $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. Since $E_{i}$ is symmetric for all $i\in I$, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}=\left\{i\in I\mid\left(y(i),x(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
To see that ${\leqslant_{Y}}\subseteq E_{Y}$, let $[x]\leqslant_{Y}[y]$. Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$. Since ${\leqslant_{i}}\subseteq E_{i}$ for all $i\in I$, it follows that $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and so $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. Hence, $\left([x],[y]\right)\in E_{Y}$.
The first part of item (iii) is proved as follows:
$[x]\leqslant_{Y}[y]$ iff $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha_{X}\left(x\right)\right)(i)\leqslant_{i}\left(\alpha_{X}\left(y\right)\right)(i)\right\}\in\mathcal{F}$
iff $\left[\alpha_{X}(x)\right]\leqslant_{Y}\left[\alpha_{X}(y)\right]$
iff $\alpha_{Y}\left([x]\right)\leqslant_{Y}\alpha_{Y}\left([y]\right).$
The second equivalence above follows from the equality
$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}=\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\},$
which is a consequence of the fact that $\alpha_{i}$ is an order automorphism of $\mathbf{X}_{i}$ for all $i\in I$.
To show that $\alpha_{Y}\subseteq E_{Y}$, let $\left([x],[y]\right)\in\alpha_{Y}$. Then $\alpha_{Y}\left([x]\right)=[y]$, and so $\left[\alpha_{X}(x)\right]=[y]$. Hence, $\left(\alpha_{X}(x),y\right)\in E_{\mathcal{F}}$, and so $\left\{i\in I\mid\left(\alpha_{X}(x)\right)(i)=y(i)\right\}=\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\in\mathcal{F}$. Since $\alpha_{i}\subseteq E_{i}$ for all $i\in I$, we have $\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and therefore, since $\mathcal{F}$ is closed under taking supersets, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$. We thus get $\left([x],[y]\right)\in E_{Y}$.
Item (iv) follows from part (i) of Theorem 2.7.
Before proving that the class $\mathsf{RDInFL}$ is closed under ultraproducts, we need the lemmas below to ensure that we have a candidate for the injective homomorphism needed in Theorem 4.11. For Lemmas 4.7 and 4.9, and Theorem 4.11, we will be working in the following setting. Let $I$ be an index set and $\mathcal{F}$ an ultrafilter on $I$. For each $i\in I$, let $\mathbf{A}_{i}$ be a representable distributive InFL-algebra. Let $\mathbf{A}$ denote the direct product $\prod\left\{\mathbf{A}_{i}\mid i\in I\right\}$, and $\mathbf{A}/\theta_{\mathcal{F}}$ the ultraproduct.
We know that for each $i\in I$, there is a poset $\left\langle X_{i},\leqslant_{i}\right\rangle$, an equivalence relation $E_{i}$ on $X_{i}$ such that ${\leqslant_{i}}\subseteq E_{i}$, an order automorphism $\alpha_{i}$ of $\left\langle X_{i},\leqslant_{i}\right\rangle$ such that $\alpha_{i}\subseteq E_{i}$, and an embedding $\varphi_{i}$ from $\mathbf{A}_{i}$ into $\mathfrak{D}\left(\left\langle E_{i},\preccurlyeq_{i}\right\rangle\right)$. We may assume that $X_{i}\neq\varnothing$ for all $i\in I$ since the direct product of a family of algebras from $\mathsf{EDInFL}$ is isomorphic to another such product from which all factors of the form $\mathfrak{D}\left(\left\langle\varnothing,\varnothing\right\rangle\right)$ have been deleted. Now, for each $[a]\in A/\theta_{\mathcal{F}}$, define
$\left([x],[y]\right)\in R_{[a]}\qquad\textnormal{iff}\qquad\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}.$
Lemma 4.7
For any $[a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
Proof 4.8
Let $\left([x],[y]\right)\in R_{[a]}$ and assume $\left([x],[y]\right)\preccurlyeq_{Y}\left([u],[v]\right)$. The first statement implies that we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$. From the second part we get $[u]\leqslant_{Y}[x]$ and $[y]\leqslant_{Y}[v]$, which means $\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}$. Since $\mathcal{F}$ is closed under taking intersections, we have
$\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}.$
If we can show that $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$ is contained in $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}$, we can use the fact that $\mathcal{F}$ is closed under taking supersets to conclude that $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$, which means $\left([u],[v]\right)\in R_{[a]}$. To this end, let $j\in\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$. Then $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$, $u(j)\leqslant_{j}x(j)$ and $y(j)\leqslant_{j}v(j)$. From the second and third part we obtain $\left(x(j),y(j)\right)\preccurlyeq_{j}\left(u(j),v(j)\right)$. Hence, since $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$ and $\varphi_{j}\left(a(j)\right)$ is an upset of $\left\langle E_{j},\preccurlyeq_{j}\right\rangle$, we have $\left(u(j),v(j)\right)\in\varphi_{j}\left(a(j)\right)$.
With the above lemma in hand, we can define a function $\varphi:A/\theta_{\mathcal{F}}\to\mathsf{Up}\left(\langle E_{Y},\preccurlyeq_{Y}\rangle\right)$ by setting $\varphi\left([a]\right)=R_{[a]}$ for all $\left[a\right]\in A/\theta_{\mathcal{F}}$.
Lemma 4.9
The map $\varphi$ is an injective function from $A/\theta_{\mathcal{F}}$ to $\mathsf{Up}\left(\langle E_{Y},\preccurlyeq_{Y}\rangle\right)$.
Proof 4.10
By Lemma 4.7 we know that $\varphi$ is well-defined. Now assume that $[a]\neq[b]$. This implies $(a,b)\notin\theta_{\mathcal{F}}$ and hence $\{i\in I\mid a(i)=b(i)\}\notin\mathcal{F}$. We get $\{i\in I\mid a(i)\neq b(i)\}\in\mathcal{F}$, and since each of the $\varphi_{i}$ is injective, it follows that $\{i\in I\mid a(i)\neq b(i)\}\subseteq\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}$. But $\mathcal{F}$ is closed under taking supersets, so $\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}\in\mathcal{F}$. Now
$\{i\in I\mid\varphi_{i}(a(i))\neq\varphi_{i}(b(i))\}=\{i\in I\mid\varphi_{i}(a(i))\nsubseteq\varphi_{i}(b(i))\}\cup\{i\in I\mid\varphi_{i}(a(i))\nsupseteq\varphi_{i}(b(i))\}\in\mathcal{F}.$
Hence, either $\{i\in I\mid\varphi_{i}(a(i))\nsubseteq\varphi_{i}(b(i))\}\in\mathcal{F}$ or $\{i\in I\mid\varphi_{i}(a(i))\nsupseteq\varphi_{i}(b(i))\}\in\mathcal{F}$, since $\mathcal{F}$ is an ultrafilter. Let $V=\{i\in I\mid\varphi_{i}(a(i))\nsubseteq\varphi_{i}(b(i))\}$ and assume $V\in\mathcal{F}$. Since $\mathcal{F}$ is a proper filter, $V\neq\varnothing$. For each $i\in V$, there exists $(s_{i},t_{i})\in E_{i}$ such that $(s_{i},t_{i})\in\varphi_{i}(a(i))$ and $(s_{i},t_{i})\notin\varphi_{i}(b(i))$. Since each $X_{i}\neq\varnothing$, we have $r_{i}\in X_{i}$ for all $i\in I$. Now define two elements of $\prod_{i\in I}X_{i}$ as follows:
$x(i)=\begin{cases}s_{i}&i\in V\\ r_{i}&i\notin V\end{cases}\qquad\text{ and }\qquad y(i)=\begin{cases}t_{i}&i\in V\\ r_{i}&i\notin V.\end{cases}$
It is not difficult to see that
$V\subseteq\{i\in I\mid(x(i),y(i))\in\varphi_{i}(a(i))\}\quad\text{and}\quad V\subseteq\{i\in I\mid(x(i),y(i))\notin\varphi_{i}(b(i))\},$
and hence $\{i\in I\mid(x(i),y(i))\in\varphi_{i}(a(i))\}\in\mathcal{F}$ and $\{i\in I\mid(x(i),y(i))\notin\varphi_{i}(b(i))\}\in\mathcal{F}$. The second part implies that $\{i\in I\mid(x(i),y(i))\in\varphi_{i}(b(i))\}^{c}\in\mathcal{F}$, which means
$\{i\in I\mid(x(i),y(i))\in\varphi_{i}(b(i))\}\notin\mathcal{F}.$
Hence we have $([x],[y])\in\varphi([a])$ but $([x],[y])\notin\varphi([b])$. That is, $\varphi([a])\neq\varphi([b])$. The proof of the case $\{i\in I\mid\varphi_{i}(a(i))\nsupseteq\varphi_{i}(b(i))\}\in\mathcal{F}$ is similar.
The function $\varphi$ from Lemma 4.9 is now shown to be a homomorphism.
Theorem 4.11
$\mathbb{P}_{\mathbb{U}}\left(\mathsf{RDInFL}\right)=\mathsf{RDInFL}$.
Proof 4.12
We have to show that the ultraproduct $\mathbf{A}/\theta_{\mathcal{F}}$ of $\left\{\mathbf{A}_{i}\mid i\in I\right\}$ is representable. Recall that each of the $\mathbf{A}_{i}$ is embedded via $\varphi_{i}$ into an algebra of the form $\mathfrak{D}\left(\langle E_{i},\preccurlyeq_{i}\rangle\right)$. Each $\langle E_{i},\preccurlyeq_{i}\rangle$ comes from a poset $\mathbf{X}_{i}$ equipped with the necessary $E_{i}$ and $\alpha_{i}$. The definitions after Theorem 4.3 describe the structure $\langle E_{Y},\preccurlyeq_{Y}\rangle$, and, by Theorem 4.5(iv), $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},\sim,-\right\rangle$ is a distributive InFL-algebra.
From Lemma 4.9 we have that $\varphi$ is injective. We will now show that $\varphi$ is a homomorphism from $\mathbf{A}/\theta_{\mathcal{F}}$ into $\mathfrak{D}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right)$. We first show that $\varphi$ preserves meets:
$\left([x],[y]\right)\in\varphi\left([a]\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)$
iff $\left([x],[y]\right)\in\varphi\left(\left[a\wedge^{\mathbf{A}}b\right]\right)$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\wedge^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\wedge^{\mathbf{A}_{i}}b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\cap\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cap\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\textnormal{ and }\left([x],[y]\right)\in\varphi\left([b]\right)$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\cap\varphi\left([b]\right).$
The fourth equivalence above follows from the fact that $\varphi_{i}$ preserves meets for all $i\in I$. The left-to-right implication of the the third-to-last equivalence follows from the fact that $\mathcal{F}$ is closed under taking supersets, and the right-to-left implication follows from the fact that $\mathcal{F}$ is closed under taking intersections.
Using the facts that $\mathcal{F}$ is an ultrafilter and $\varphi_{i}$ preserves joins for all $i\in I$, we can also show that $\varphi$ preserves joins:
$\left([x],[y]\right)\in\varphi\left([a]\vee^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)$
iff $\left([x],[y]\right)\in\varphi\left(\left[a\vee^{\mathbf{A}}b\right]\right)$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\vee^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\vee^{\mathbf{A}_{i}}b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\cup\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\textnormal{ or }\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cup\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}\textnormal{ or }\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\textnormal{ or }\left([x],[y]\right)\in\varphi\left([b]\right)$
iff $\left([x],[y]\right)\in\varphi\left([a]\right)\cup\varphi\left([b]\right).$
To show that $\varphi$ preserves the monoid operation, first let $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$. Then there is some $[z]\in Y$ such that $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$. Hence, we obtain $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$, which means
$\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cap\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}.$
Since $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cap\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}$ is contained in the set $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\circ\varphi_{i}\left(b(i)\right)\right\}$ and $\mathcal{F}$ is closed under taking supersets, it follows that $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\circ\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$. Now $\varphi_{i}\left(a(i)\right)\circ\varphi_{i}\left(b(i)\right)=\varphi_{i}\left(a(i)\cdot^{\mathbf{A}_{i}}b(i)\right)=\varphi_{i}\left(\left(a\cdot^{\mathbf{A}}b\right)(i)\right)$, so we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\cdot^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F}$. This shows that $\left([x],[y]\right)\in\varphi\left(\left[a\cdot^{\mathbf{A}}b\right]\right)=\varphi\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)$.
For the reverse inclusion, let $\left([x],[y]\right)\in\varphi\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)=\varphi\left(\left[a\cdot^{\mathbf{A}}b\right]\right)$. Then we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\cdot^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F},$ and so
$\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\circ\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}.$
Let $V=\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\circ\varphi_{i}\left(b(i)\right)\right\}$. Since $\mathcal{F}$ is a proper filter, $V\neq\varnothing$. For each $i\in V$, there exists some $s_{i}\in X_{i}$ such that $\left(x(i),s_{i}\right)\in\varphi_{i}\left(a(i)\right)$ and $\left(s_{i},y(i)\right)\in\varphi_{i}\left(b(i)\right)$. Now define an element of $\prod_{i\in I}X_{i}$ as follows:
$z(i)=\begin{cases}s_{i}&i\in V\\ x(i)&i\notin V\end{cases}$
It then follows that $V\subseteq\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cap\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}$ Hence, since $\mathcal{F}$ is closed under taking supersets, we have
$\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cap\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
which means $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$. This shows $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$, and therefore $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$.
Next, we show that $\varphi$ preserves the monoid identity:
$\left([x],[y]\right)\in\varphi\left(1^{\mathbf{A}/\theta_{\mathcal{F}}}\right)$ iff $\left([x],[y]\right)\in\varphi\left(\left[1^{\mathbf{A}}\right]\right)$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(1^{\mathbf{A}}(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(1^{\mathbf{A}_{i}}\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$
iff $[x]\leqslant_{Y}[y].$
We now prove that $\varphi$ preserves --:
$\left([x],[y]\right)\in\varphi\left(-^{\mathbf{A}/\theta_{\mathcal{F}}}[a]\right)=\varphi\left(\left[-^{\mathbf{A}}a\right]\right)$ iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(-^{\mathbf{A}}a\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(-^{\mathbf{A}_{i}}a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in-\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\alpha_{i}\circ\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha_{i}\left(x(i)\right),y(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}.$
We recall that the complement $\left(\varphi_{i}\left(a(i)\right)\right)^{c}$ is being taken in $E_{i}$, hence
$\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\textnormal{ and }\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\cap\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}^{c}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left([y],[\alpha_{X}(x)]\right)\in E_{Y}\textnormal{ and }\left([y],[\alpha_{X}(x)]\right)\notin\varphi\left([a]\right)$
iff $\left([y],[\alpha_{X}(x)]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
iff $\left([\alpha_{X}(x)],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left(\alpha_{Y}\left([x]\right),[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left([x],[y]\right)\in\alpha_{Y}\circ\left(\varphi\left([a]\right)\right)^{c\smile}=-\varphi\left([a]\right).$
Using Lemma 4.4, we can also show that $\varphi$ preserves $\sim$:
$\left([x],[y]\right)\in\varphi\left({\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}[a]\right)=\varphi\left(\left[{\sim}^{\mathbf{A}}a\right]\right)$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left({\sim}^{\mathbf{A}}a\right)(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left({\sim}^{\mathbf{A}_{i}}a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in{\sim}\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\circ\alpha_{i}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(x(i),\alpha^{-1}_{i}\left(y(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}.$
Again recalling that the complement $\left(\varphi_{i}\left(a(i)\right)\right)^{c}$ is taken in $E_{i}$, we get:
$\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in E_{i}\textnormal{ and }\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in E_{i}\right\}\cap\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}^{c}\in\mathcal{F}$
iff $\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(\alpha^{-1}_{i}\left(y(i)\right),x(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left\{i\in I\mid\left(\left(\alpha^{-1}_{X}\left(y\right)\right)(i),x(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(\left(\alpha^{-1}_{X}\left(y\right)\right)(i),x(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
iff $\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right).$
Using the fact that the complement $\left(\varphi([a])\right)^{c}$ is taken in $E_{Y}$, we have
$\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right)$
iff $\left([\alpha^{-1}_{X}(y)],[x]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
iff $\left([x],[\alpha^{-1}_{X}(y)]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left([x],\alpha^{-1}_{Y}\left([y]\right)\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
iff $\left([x],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}\circ\alpha_{Y}={\sim}\varphi\left([a]\right).$
Theorem 4.11 finds an immediate application in this final subsection. By combining it with Definition 4.1, we get the following result:
Corollary 4.13
$\mathbb{P_{U}}(\mathsf{RSM})=\mathsf{RSM}$.
Proof 4.14
Consider $\{\,\mathbf{A}_{i}\mid i\in I\,\}\subseteq\mathsf{RSM}$. Each of the $\mathbf{A}_{i}$ is isomorphic to a direct reduct $\left(\mathbf{B}_{i}\right)_{r}$ of an algebra $\mathbf{B}_{i}\in\mathsf{RDInFL}$. That is, for each $i\in I$, there is a map $\varphi_{i}:A_{i}\to B_{i}$ which is an isomorphism from $\mathbf{A}_{i}$ to $\left(\mathbf{B}_{i}\right)_{r}$.
Let $\mathcal{F}$ be an ultrafilter on $I$ and recall that $\mathbf{A}$ and $\mathbf{B}$ are the direct products of the $\mathbf{A}_{i}$ and $\mathbf{B}_{i}$, respectively. (We denote by $A$ and $B$ the underlying sets of $\mathbf{A}$ and $\mathbf{B}$ respectively.) By Theorem 4.11, we have that $\mathbf{B}/\theta_{\mathcal{F}}\in\mathsf{RDInFL}$. We will show that $\mathbf{A}/\theta_{\mathcal{F}}$ is isomorphic to $\left(\mathbf{B}/\theta_{\mathcal{F}}\right)_{r}$.
Let $\varphi:A\to B$ be defined by $\varphi(a)(i)=\varphi_{i}(a(i))$. Clearly $\varphi$ is a bijective map. Next, consider $\overline{\varphi}:A/\theta_{\mathcal{F}}\to B/\theta_{\mathcal{F}}$, defined by $\overline{\varphi}([a])=[\varphi(a)]$. We show that $\overline{\varphi}$ is bijective. For $[b]\in B/\theta_{\mathcal{F}}$, since $b\in B$, there exists $a\in A$ with $\varphi(a)=b$. So $\overline{\varphi}\left([a]\right)=\left[\varphi(a)\right]=[b]$. If $a,c\in A$ with $[a]\neq[c]$, then $(a,c)\notin\theta_{\mathcal{F}}$. Thus $\{\,i\in I\mid a(i)=c(i)\,\}\notin\mathcal{F}$, so $\{\,i\in I\mid a(i)\neq c(i)\,\}\in\mathcal{F}$. Since the $\varphi_{i}$ are injective, if $a(i)\neq c(i)$, then $\varphi_{i}(a(i))\neq\varphi_{i}(c(i))$ and so $\varphi(a)(i)\neq\varphi(c)(i)$. Hence $\{\,i\in I\mid a(i)\neq c(i)\,\}\subseteq\{\,i\in I\mid\varphi(a)(i)\neq\varphi(c)(i)\,\}$ so the latter set is in $\mathcal{F}$. Now $\{\,i\in I\mid\varphi(a)(i)=\varphi(c)(i)\,\}\notin\mathcal{F}$, so $(\varphi(a),\varphi(c))\notin\theta_{\mathcal{F}}$ and therefore $[\varphi(a)]\neq[\varphi(c)]$, i.e. $\overline{\varphi}\left([a]\right)\neq\overline{\varphi}\left([c]\right)$.
Next, we will show that $\overline{\varphi}$ preserves the fusion operation. For $[a],[c]\in A/\theta_{\mathcal{F}}$, we have
$\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left(\left[a\cdot^{\mathbf{A}}c\right]\right)=\left[\varphi\left(a\cdot^{\mathbf{A}}c\right)\right].$
For each $i\in I$, $\varphi_{i}:\mathbf{A}_{i}\to(\mathbf{B}_{i})_{r}$ is an isomorphism. Also, $\cdot^{(\mathbf{B}_{i})_{r}}=\cdot^{\mathbf{B}_{i}}$. Hence for any $i\in I$:
$\varphi(a\cdot^{\mathbf{A}}c)(i)=\varphi_{i}\left((a\cdot^{\mathbf{A}}c)(i)\right)=\varphi_{i}\left(a(i)\cdot^{\mathbf{A}_{i}}c(i)\right)=\varphi_{i}(a(i))\cdot^{(\mathbf{B}_{i})_{r}}\varphi_{i}(c(i))=\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i)).$
Also observe that $\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i))=\left(\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right)(i)$. Using this in the last step below, we can now see that
$\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)=\overline{\varphi}\left([a]\right)\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\overline{\varphi}\left([c]\right)=[\varphi(a)]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}[\varphi(c)]=\left[\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right]=\left[\varphi(a\cdot^{\mathbf{A}}c)\right]$
and hence $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)$. Using the same approach, it is straightforward to show that $\overline{\varphi}$ preserves the operations $\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}$, $\vee^{\mathbf{A}/\theta_{\mathcal{F}}}$, $1^{\mathbf{A}/\theta_{\mathcal{F}}}$ and ${\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}$.
Now consider the operation $\to^{\mathbf{A}/\theta_{\mathcal{F}}}$ on $\mathbf{A}/\theta_{\mathcal{F}}$. For $[a]$, $[c]\in A/\theta_{\mathcal{F}}$, we have
$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)=\left[\varphi(a\to^{\mathbf{A}}c)\right].$
For $i\in I$ and $b,d\in B_{i}$, recall that $b\to^{(\mathbf{B}_{i})_{r}}d={\sim}^{\mathbf{B}_{i}}\left(-^{\mathbf{B}_{i}}d\cdot^{\mathbf{B}_{i}}b\right)$. The third equality below uses the fact that $\varphi_{i}:\mathbf{A}_{i}\to(\mathbf{B}_{i})_{r}$ is an isomorphism. For any $i\in I$, we get:
$\varphi(a\to^{\mathbf{A}}c)(i)=\varphi_{i}\left((a\to^{\mathbf{A}}c)(i)\right)$ $=\varphi_{i}\left(a(i)\to^{\mathbf{A}_{i}}c(i)\right)$
$=\varphi_{i}(a(i))\to^{(\mathbf{B}_{i})_{r}}\varphi_{i}(c(i))$
$={\sim}^{\mathbf{B}_{i}}\left(-^{\mathbf{B}_{i}}\varphi_{i}(c(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(a(i))\right)$
$={\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)(i).$
Finally, we get:
$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)$ $=\left[\varphi(a\to^{\mathbf{A}}c)\right]$
$=\left[{\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
$={\sim}^{\mathbf{B}/\theta_{\mathcal{F}}}\left[\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
$={\sim}^{\mathbf{B}/\theta_{\mathcal{F}}}\left(\big{[}-^{\mathbf{B}}\varphi(c)\big{]}\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\big{[}\varphi(a)\big{]}\right)$
$={\sim}^{\mathbf{B}/\theta_{\mathcal{F}}}\left(-^{\mathbf{B}/\theta_{\mathcal{F}}}\left[\varphi(c)\right]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\left[\varphi(a)\right]\right)$
$=\left[\varphi(a)\right]\to^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\left[\varphi(c)\right]$
$=\overline{\varphi}\left([a]\right)\to^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right).$
4.2 Representing all Sugihara monoids
In addition to $\mathsf{RSM}$ being closed under ultraproducts, it follows from the definition of $\mathsf{RDInFL}$ that $\mathsf{RSM}$ is closed under $\mathbb{I}$, $\mathbb{S}$ and $\mathbb{P}$. Next, we recall the following well-known theorem from universal algebra (cf. [2, Theorem V-2.14]).
Theorem 4.15
Every algebra can be embedded in an ultraproduct of its finitely generated subalgebras.
The finitely generated subalgebras of $\mathbf{S}=(\mathbb{Z},\wedge,\vee,\cdot,\to,1,{\sim})$ are the algebras $\mathbf{S}_{2n+1}$ for $n\in\omega$. For the infinite Sugihara monoid $\mathbf{S}^{*}=(\mathbb{Z}{\setminus}\{0\},\wedge,\vee,\cdot,\to,1,{\sim})$, its finitely generated subalgebras are $\mathbf{S}_{2n+2}$ for $n\in\omega$. Lastly, recall that the quasivariety of Sugihara monoids is generated by $\{\mathbf{S},\mathbf{S}^{*}\}$ (cf. [17, Section 3]). We now have everything in place in order to prove our main result.
Theorem 4.16
Every Sugihara monoid is representable. That is, $\mathsf{SM}\subseteq\mathsf{RSM}$.
Proof 4.17
We need to show that $\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})=\mathsf{SM}$ is contained in $\mathsf{RSM}$. From Theorem 3.22 and Examples 2.92.10 and 2.11, we know that $\{\mathbf{S}_{i}\mid i\in\mathbb{Z}^{+}\}\subseteq\mathsf{RSM}$. (The trivial algebra $\mathbf{S}_{1}$ is of course representable as it is a subalgebra of $\mathbf{S}_{3}$.)
[…]
Since $\mathsf{SM}=\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})$ and $\{\mathbf{S},\mathbf{S}^{*}\}\subseteq\mathsf{RSM}$ we get $\mathsf{SM}=\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})\subseteq\mathbb{ISP}(\mathsf{RSM})=\mathsf{RSM}$.
Acknowledgements
The authors are extremely grateful to the two anonymous referees for their very insightful and knowledgeable feedback on the first version on this paper. Their comments enabled us to see that additional results could be possible, and this led to the development of Section 4. The first author would like to thank Chapman University for its hospitality during a research visit in September/October 2023. We would also like to thank Jamie Wannenburg for useful discussions on this topic.

### v3

@@
 The first step is to obtain an explicit relational representation of all finite odd Sugihara chains.
-Our construction mimics that of Maddux (2010), where a relational representation of the finite even Sugihara chains is given.
+Our construction mimics that of Maddux [16], where a relational representation of the finite even Sugihara chains is given.
 We define the class of representable Sugihara monoids as those which can be represented as so-called direct reducts of distributive involutive FL-algebras of binary relations.
@@
 This map $\alpha$ is an order isomorphism of the poset consisting of two chains.
-The algebra $\mathbf{S}_{2n+3}$ is then embedded in the upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$, where this time $\mathbf{X}$ consists of two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$.
+The algebra $\mathbf{S}_{2n+3}$ is then embedded in the upsets of $\mathbf{X}^{\partial}\times\mathbf{X}$, where this time $\mathbf{X}$ consists of two disjoint copies of $\langle\mathbb{Q}^{n},\leqslant_{n}\rangle$ with order relations between them (see Figure 1).
 The key step is that $\alpha$ is used to define the order-reversing involution in a way that ensures the algebra will be an odd Sugihara monoid.
@@
 It does so by taking the complement inside a symmetric relation, which does not have to be $X^{2}$.
-The map $\alpha$ also plays a crucial role in the representation of distributive involutive FL-algebras.
-These algebras have two order-reversing maps $\sim$ and -- (see Sections 2.1 and 2.4).
+We will view Sugihara monoids as examples of distributive involutive FL-algebras (DInFL-algebras).
+See Section 2.1 for a precise definition of these algebras.
+DInFL-algebras have two order-reversing operations $\sim$ and --.
 If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic.
@@
 This gives us algebras $\mathbf{S}_{n}=\langle S_{n},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$.
-(We remark that when we talk of a Sugihara chain, we mean an algebra that is a Sugihara monoid and is linearly ordered.
-This is different to the terminology used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
+(When we talk of a Sugihara chain, we mean a linearly ordered Sugihara monoid, using slightly different terminology from that used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
 […]
@@
 If $E=X^{2}$ then we say that $\mathfrak{D}(\mathbf{E})$ is a full distributive involutive FL-algebra, and denote the class of such algebras by $\mathsf{FDInFL}$.
-Before we give some examples, we need the following definition. It is an adaptation of [14, Definition 8].
+Before we give some examples, we need the definition of a direct reduct.
+This is an adaptation of [14, Definition 8].
+It differs from an ordinary reduct because the set of operations is not simply a subset of the original, but new operations are also defined in terms of the old operations.
 Definition 2.8
@@
 Example 2.9
-Consider the two-element Sugihara chain $\mathbf{S}_{2}$.
-The underlying set is given by $S_{2}=\{a_{-1},a_{1}\}$ where $a_{1}$ is the identity of the monoid operation.
-Consider the one-element poset with $X=\{x\}$ and order ${\leqslant}=\mathrm{id}_{X}$.
-Let $E=X^{2}$ and $\alpha=\mathrm{id}_{X}$.
-Then $\left\langle\mathsf{Up}\left(\left\langle X^{2},\leqslant\right\rangle\right),\cap,\cup,\circ,\leqslant,\sim,-\right\rangle$ is a two-element distributive InFL-algebra whose direct reduct is isomorphic to $\mathbf{S}_{2}$.
+The two-element Sugihara chain $\mathbf{S}_{2}$ can be represented over $X=\{x\}$ with ${\leqslant}=E=\alpha=\mathrm{id}_{X}$.
 Using Theorem 2.7, we obtain a relational representation of the odd Sugihara chain $\mathbf{S}_{3}$. We observe that this representation is not covered by the construction given in Section 3.
@@
     for the DInFL-algebras $\mathfrak{D}\left(\mathbf{E}\right)$ and $\prod_{i\in I}\mathfrak{D}\left(\mathbf{E}_{i}\right)$, we have $\mathfrak{D}\left(\mathbf{E}\right)\cong\displaystyle\prod_{i\in I}\mathfrak{D}\left(\mathbf{E}_{i}\right)$.
-The following theorem relies on part (iii) of Theorem 2.12 and is used to representable DInFL-algebras.
+The following theorem relies on part (iii) of Theorem 2.12 and is used to define representable DInFL-algebras.
 Theorem 2.13
-[3, Theorem 4.4] $\mathbb{P}\left(\mathsf{FDInFL}\right)=\mathsf{EDInFL}$
+[3, Theorem 4.4] $\mathbb{P}\left(\mathsf{FDInFL}\right)=\mathsf{EDInFL}$.
 The definition below is a generalisation of [3, Definition 4.5], given there for DqRAs.
@@
 A DInFL-algebra $\mathbf{A}=\left\langle A,\wedge,\vee,\cdot,1,{\sim},{-}\right\rangle$ is representable if $\mathbf{A}\in\mathbb{ISP}\left(\mathsf{FDInFL}\right)$ or, equivalently, $\mathbf{A}\in\mathbb{IS}\left(\mathsf{EDInFL}\right)$.
-In Theorem 4.11 we will show that the class of representable DInFL-algebras is closed under ultraproducts.
+We denote the class of representable DInFL-algebras by $\mathsf{RDInFL}$.
+In Theorem 4.12 we will show that $\mathsf{RDInFL}$ is closed under ultraproducts.
 That will be a key step in showing that all Sugihara monoids can be represented as algebras of binary relations.
@@
 Then $x^{b}=y^{d}$ or $x<y$.
-If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$.
-Hence, $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, i.e., $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}\circ\alpha=0$.
+If $x^{b}=y^{d}$, then $y^{-d}\not\leqslant x^{b}$, and so $\left(x^{b},y^{d}\right)\in 0$.
 On the other hand, if $x<y$, then $y\not\leqslant x$, so $y^{-d}\not\leqslant x^{b}$.
-Again we get $\left(x^{b},y^{d}\right)\in\left(\leqslant_{X}\right)^{c\smile}=0$.
+Again we have $(x^{b},y^{d})\in 0$.
 For the other inclusion, let $\left(x^{b},y^{d}\right)\in 0=\left(\leqslant_{X}\right)^{c\smile}\circ\alpha$.
-Then we have $\left(x^{b},\alpha(y^{d})\right)\in{\leqslant^{c\smile}}$, which means $\left(x^{b},y^{-d}\right)\in{\leqslant^{c\smile}}$.
-We therefore get $y^{-d}\not\leqslant x^{b}$.
-If $b=-d$, then it must be the case that $y\not\leqslant x$, in which case we have $x<y$.
+Then we have $y^{-d}\not\leqslant x^{b}$.
+If $b=-d$, then $y\not\leqslant x$, so $x<y$.
 Hence, $x^{b}\leqslant_{X}y^{d}$.
@@
 Now set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$.
-The lemma below will help to simplify some of the proofs later on.
+Then $\leqslant_{n}$ is the lexicographic order on $\mathbb{Q}^{n}$.
+Since the rational numbers is linearly ordered, it follows that $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
+Part (i) of the lemma below can be used to show that $\leqslant_{n}$ is transitive.
+It will also help to simplify some of the proofs later on.
 Lemma 3.3
-Let $n\geqslant 1$ and $j,k\in\left\{1,\ldots,n\right\}$. If $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$, then $\left(p,r\right)\in L_{\textnormal{min}\{j,k\}}$.
-Proof 3.4
-Let $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$.
-Then $\left(p_{1},\ldots,p_{j-1}\right)=\left(q_{1},\ldots,q_{j-1}\right)$, $p_{j}<_{\mathbb{Q}}q_{j}$, $\left(q_{1},\ldots,q_{k-1}\right)=\left(r_{1},\ldots,r_{k-1}\right)$ and $q_{k}<_{\mathbb{Q}}r_{k}$.
-Clearly then $\left(p_{1},\ldots,p_{\textnormal{min}\{j,k\}-1}\right)=\left(r_{1},\ldots,r_{\textnormal{min}\{j,k\}-1}\right)$.
-If $j=k$, then we get $p_{j}<q_{j}<r_{j}$.
-If $\text{min}\{j,k\}=j$, then $p_{j}<q_{j}=r_{j}$.
-Finally, if $\text{min}\{j,k\}=k$, then $p_{k}=p_{k}=q_{k}<r_{k}$.
-Hence, in all three cases we have $(p,r)\in L_{\text{min}\{j,k\}}$.
-Lemma 3.5
-Let $n\geqslant 1$. Then $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
-Proof 3.6
-The relation $\leqslant_{n}$ is clearly reflexive and anti-symmetric.
-For transitivity, assume $p\leqslant_{n}q$ and $q\leqslant_{n}r$.
-If $p=q$ or $q=r$, then we clearly have $p\leqslant_{n}r$.
-Assume there are $j,k\in\left\{1,\ldots,n\right\}$ such that $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$.
-In this case we have $\left(p,r\right)\in L_{{\textnormal{min}\{j,k\}}}$ by Lemma 3.3, which means $p\leqslant_{n}r$.
-Finally, we show that $\leqslant_{n}$ is a linear order.
-If $p=q$, we are done, so suppose $p\neq q$.
-Then there is a smallest $i\in\left\{1,\ldots,n\right\}$ such that $p_{i}\neq q_{i}$.
-Since the set of rational numbers is linearly ordered, we have $p_{i}<_{\mathbb{Q}}q_{i}$ or $q_{i}<_{\mathbb{Q}}p_{i}$.
-If $i=1$, then $\left(p,q\right)\in L_{1}$ or $\left(q,p\right)\in L_{1}$, which means $p\leqslant_{n}q$ or $q\leqslant_{n}p$.
-On the other hand, if $i\in\{2,\ldots,n\}$, then $p_{j}=q_{j}$ for all $j\in\left\{1,\ldots,i-1\right\}$.
-Hence, $\left(p,q\right)\in L_{i}$ or $\left(q,p\right)\in L_{i}$, and therefore $p\leqslant_{n}q$ or $q\leqslant_{n}p$.
+Let $n\geqslant 1$.
+  1. (i)
+    If $\left(p,q\right)\in L_{j}$ and $\left(q,r\right)\in L_{k}$ with $j,k\in\{1,\ldots,n\}$, then $\left(p,r\right)\in L_{\textnormal{min}\{j,k\}}$.
+  2. (ii)
+    The structure $\left\langle\mathbb{Q}^{n},\leqslant_{n}\right\rangle$ is a chain.
 Let $X:=\left(\mathbb{Q}^{n}\times\left\{-1\right\}\right)\cup\left(\mathbb{Q}^{n}\times\left\{1\right\}\right)$, and define a binary relation $\leqslant_{X}$ on $X$ by
@@
 For each $i\in\left\{-n-1,\ldots,n+1\right\}$, the relation $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$.
-Proof 3.8
+Proof 3.5
 Both $X^{2}$ and $\varnothing$ are upsets of $\left\langle X^{2},\preccurlyeq\right\rangle$.
@@
 For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $-R_{i}={\sim}R_{i}$.
-Proof 3.10
+Proof 3.7
 If $i=0$, the claim follows from Lemma 2.5. Now consider the following cases:
@@
 $R_{-n-1}\subseteq R_{-n}\subseteq\cdots\subseteq R_{-1}\subseteq R_{0}\subseteq R_{1}\subseteq\cdots\subseteq R_{n}\subseteq R_{n+1}.$
-Proof 3.12
+Proof 3.9
 It follows from the construction that $R_{-i}\subseteq R_{-i+1}$ for all $i\in\left\{2,\ldots,n+1\right\}$.
@@
 For each $i\in\left\{-n-1,\ldots,n+1\right\}$, the relation $R_{i}$ is transitive.
-Proof 3.14
+Proof 3.11
 This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for $i=-n-1$. We now consider the following two cases:
@@
 If $|i|>|j|$ then $R_{i}\circ R_{j}=R_{i}$.
-Proof 3.16
+Proof 3.13
 If $i=-n-1$ (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
@@
 If $|j|>|i|$ then $R_{i}\circ R_{j}=R_{j}$.
-Proof 3.18
+Proof 3.15
 If $j=-n-1$, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
@@
 If $|j|=|i|$ then $R_{i}\circ R_{j}=R_{\text{min}\{i,j\}}$.
-Proof 3.20
+Proof 3.17
 The claim clearly follows for the cases $i=j=0$ and $i=j=-n-1$. So consider the following cases:
@@
 Let $n\geqslant 1$. Then the algebra $\mathbb{S}_{2n+3}$ is an odd Sugihara chain and $\mathbf{S}_{2n+3}\cong\mathbb{S}_{2n+3}$.
-Proof 3.23
+Proof 3.20
 […]
@@
 An ultraproduct of a family of DInFL-algebras is again a DInFL-algebra.
+Below we outline how to construct the ultraproduct of a set of partially ordered sets with additional structure, each of which can be used to represent a DInFL-algebra.
 Consider an index set $I$ and let $\mathcal{F}$ be an ultrafilter on $I$.
@@
 Proof 4.6
-We first prove (i).
-Let $[x]\in Y$.
-Since $\leqslant_{i}$ is reflexive for all $i\in I$, we get $\left\{i\in I\mid x(i)\leqslant_{i}x(i)\right\}=I$.
-Now $\mathcal{F}$ is an ultrafilter, so $I\in\mathcal{F}$.
-Hence, $\left[x\right]\leqslant_{Y}\left[x\right]$.
-Next, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[x]$.
-Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$.
-Since $\mathcal{F}$ is closed under taking intersections, it follows that
-$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}.$
-Using the fact that $\leqslant_{i}$ is anti-symmetric for all $i\in I$, we can show that
-$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}x(i)\right\}=\left\{i\in I\mid x(i)=y(i)\right\}.$
-Hence we get $\left\{i\in I\mid x(i)=y(i)\right\}\in\mathcal{F}$. This shows that $(x,y)\in E_{\mathcal{F}}$, and hence $[x]=[y]$.
-For the transitivity, assume $[x]\leqslant_{Y}[y]$ and $[y]\leqslant_{Y}[z]$.
-Then $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$, and so $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant z(i)\right\}\in\mathcal{F}$.
-We know that $\leqslant_{i}$ is transitive for all $i\in I$, so it follows that
-$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}z(i)\right\}\subseteq\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}.$
-Hence, since $\mathcal{F}$ is closed under taking supersets, we have $\left\{i\in I\mid x(i)\leqslant_{i}z(i)\right\}\in\mathcal{F}$, which means $[x]\leqslant_{Y}[z]$.
-We next prove item (ii).
-Firstly, since $E_{i}$ is reflexive for all $i\in I$ and $I\in\mathcal{F}$, it follows that $\left([x],[x]\right)\in E_{Y}$ for all $[x]\in Y$.
-Using a similar argument as for item (i), we can show that $E_{Y}$ is transitive.
-For symmetry, assume $\left([x],[y]\right)\in E_{Y}$.
-Then $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
-Since $E_{i}$ is symmetric for all $i\in I$, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}=\left\{i\in I\mid\left(y(i),x(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
-To see that ${\leqslant_{Y}}\subseteq E_{Y}$, let $[x]\leqslant_{Y}[y]$.
-Then we have $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$.
-Since ${\leqslant_{i}}\subseteq E_{i}$ for all $i\in I$, it follows that $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and so $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
-Hence, $\left([x],[y]\right)\in E_{Y}$.
-The first part of item (iii) is proved as follows:
-$[x]\leqslant_{Y}[y]$ iff $\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(\alpha_{X}\left(x\right)\right)(i)\leqslant_{i}\left(\alpha_{X}\left(y\right)\right)(i)\right\}\in\mathcal{F}$
-iff $\left[\alpha_{X}(x)\right]\leqslant_{Y}\left[\alpha_{X}(y)\right]$
-iff $\alpha_{Y}\left([x]\right)\leqslant_{Y}\alpha_{Y}\left([y]\right).$
-The second equivalence above follows from the equality
-$\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}=\left\{i\in I\mid\alpha_{i}\left(x(i)\right)\leqslant_{i}\alpha_{i}\left(y(i)\right)\right\},$
-which is a consequence of the fact that $\alpha_{i}$ is an order automorphism of $\mathbf{X}_{i}$ for all $i\in I$.
-To show that $\alpha_{Y}\subseteq E_{Y}$, let $\left([x],[y]\right)\in\alpha_{Y}$.
-Then $\alpha_{Y}\left([x]\right)=[y]$, and so $\left[\alpha_{X}(x)\right]=[y]$.
-Hence, $\left(\alpha_{X}(x),y\right)\in E_{\mathcal{F}}$, and so $\left\{i\in I\mid\left(\alpha_{X}(x)\right)(i)=y(i)\right\}=\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\in\mathcal{F}$.
-Since $\alpha_{i}\subseteq E_{i}$ for all $i\in I$, we have $\left\{i\in I\mid\alpha_{i}(x(i))=y(i)\right\}\subseteq\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}$, and therefore, since $\mathcal{F}$ is closed under taking supersets, we obtain $\left\{i\in I\mid\left(x(i),y(i)\right)\in E_{i}\right\}\in\mathcal{F}$.
-We thus get $\left([x],[y]\right)\in E_{Y}$.
-Item (iv) follows from part (i) of Theorem 2.7.
+Since all the conditions in items (i) to (iii) can be translated into first-order sentences in the relevant language, they follow from Łoś’s Theorem. Item (iv) follows from part (i) of Theorem 2.7.
+Remark 4.7
+The fact that the structures with which the DInFL-algebras are represented (i.e.
+posets with an equivalence relation and an order automorphism) are first-order structures, greatly simplifies the proof that $\mathsf{RDInFL}$ is closed under ultraproducts.
+Theorem 4.5 above gives us the candidate structure from which we can build the algebra in which we will embed the ultraproduct of representable algebras.
 Before proving that the class $\mathsf{RDInFL}$ is closed under ultraproducts, we need the lemmas below to ensure that we have a candidate for the injective homomorphism needed in Theorem 4.12.
-For Lemmas 4.7 and 4.9, and Theorem 4.11, we will be working in the following setting.
+For Lemmas 4.8 and 4.10, and Theorem 4.12, we will be working in the following setting.
 Let $I$ be an index set and $\mathcal{F}$ an ultrafilter on $I$.
@@
 For any $[a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
-Proof 4.8
+Proof 4.9
 Let $\left([x],[y]\right)\in R_{[a]}$ and assume $\left([x],[y]\right)\preccurlyeq_{Y}\left([u],[v]\right)$.
@@
 Since $\mathcal{F}$ is closed under taking intersections, we have
-$\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}.$
-If we can show that $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$ is contained in $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}$, we can use the fact that $\mathcal{F}$ is closed under taking supersets to conclude that $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$, which means $\left([u],[v]\right)\in R_{[a]}$.
-To this end, let $j\in\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}$.
+$J=\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\cap\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\cap\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}.$
+If we can show that $J\subseteq\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}$, we can use the fact that $\mathcal{F}$ is closed under taking supersets to conclude that $\left\{i\in I\mid\left(u(i),v(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$, which means $\left([u],[v]\right)\in R_{[a]}$.
+To this end, let $j\in J$.
 Then $\left(x(j),y(j)\right)\in\varphi_{j}(a(j))$, $u(j)\leqslant_{j}x(j)$ and $y(j)\leqslant_{j}v(j)$.
@@
 The map $\varphi$ is an injective function from $A/\theta_{\mathcal{F}}$ to $\mathsf{Up}\left(\langle E_{Y},\preccurlyeq_{Y}\rangle\right)$.
-Proof 4.10
+Proof 4.11
 By Lemma 4.8 we know that $\varphi$ is well-defined.
@@
 $\mathbb{P}_{\mathbb{U}}\left(\mathsf{RDInFL}\right)=\mathsf{RDInFL}$.
-Proof 4.12
+Proof 4.13
 We have to show that the ultraproduct $\mathbf{A}/\theta_{\mathcal{F}}$ of $\left\{\mathbf{A}_{i}\mid i\in I\right\}$ is representable.
@@
 The left-to-right implication of the the third-to-last equivalence follows from the fact that $\mathcal{F}$ is closed under taking supersets, and the right-to-left implication follows from the fact that $\mathcal{F}$ is closed under taking intersections.
-Using the facts that $\mathcal{F}$ is an ultrafilter and $\varphi_{i}$ preserves joins for all $i\in I$, we can also show that $\varphi$ preserves joins:
-$\left([x],[y]\right)\in\varphi\left([a]\vee^{\mathbf{A}/\theta_{\mathcal{F}}}[b]\right)$
-iff $\left([x],[y]\right)\in\varphi\left(\left[a\vee^{\mathbf{A}}b\right]\right)$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(a\vee^{\mathbf{A}}b\right)(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\vee^{\mathbf{A}_{i}}b(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\cup\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\textnormal{ or }\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\cup\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}\textnormal{ or }\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$
-iff $\left([x],[y]\right)\in\varphi\left([a]\right)\textnormal{ or }\left([x],[y]\right)\in\varphi\left([b]\right)$
-iff $\left([x],[y]\right)\in\varphi\left([a]\right)\cup\varphi\left([b]\right).$
+Using the facts that $\mathcal{F}$ is an ultrafilter and $\varphi_{i}$ preserves joins for all $i\in I$, we can also show that $\varphi$ preserves joins using a proof dual to the case for meets.
 To show that $\varphi$ preserves the monoid operation, first let $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$.
@@
 iff $[x]\leqslant_{Y}[y].$
-We now prove that $\varphi$ preserves --:
-$\left([x],[y]\right)\in\varphi\left(-^{\mathbf{A}/\theta_{\mathcal{F}}}[a]\right)=\varphi\left(\left[-^{\mathbf{A}}a\right]\right)$ iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(\left(-^{\mathbf{A}}a\right)(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(-^{\mathbf{A}_{i}}a(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in-\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(x(i),y(i)\right)\in\alpha_{i}\circ\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(\alpha_{i}\left(x(i)\right),y(i)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c\smile}\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}.$
-We recall that the complement $\left(\varphi_{i}\left(a(i)\right)\right)^{c}$ is being taken in $E_{i}$, hence
-$\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\left(\varphi_{i}\left(a(i)\right)\right)^{c}\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\textnormal{ and }\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\cap\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\notin\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}^{c}\in\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\alpha_{i}\left(x(i)\right)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
-iff $\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in E_{i}\right\}\in\mathcal{F}\textnormal{ and }\left\{i\in I\mid\left(y(i),\left(\alpha_{X}\left(x\right)\right)(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\notin\mathcal{F}$
-iff $\left([y],[\alpha_{X}(x)]\right)\in E_{Y}\textnormal{ and }\left([y],[\alpha_{X}(x)]\right)\notin\varphi\left([a]\right)$
-iff $\left([y],[\alpha_{X}(x)]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
-iff $\left([\alpha_{X}(x)],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
-iff $\left(\alpha_{Y}\left([x]\right),[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
-iff $\left([x],[y]\right)\in\alpha_{Y}\circ\left(\varphi\left([a]\right)\right)^{c\smile}=-\varphi\left([a]\right).$
 Using Lemma 4.4, we can also show that $\varphi$ preserves $\sim$:
@@
 Using the fact that the complement $\left(\varphi([a])\right)^{c}$ is taken in $E_{Y}$, we have
-$\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right)$
-iff $\left([\alpha^{-1}_{X}(y)],[x]\right)\in\left(\varphi\left([a]\right)\right)^{c}$
-iff $\left([x],[\alpha^{-1}_{X}(y)]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
+$\left([\alpha^{-1}_{X}(y)],[x]\right)\in E_{Y}\textnormal{ and }\left([\alpha^{-1}_{X}(y)],[x]\right)\notin\varphi\left([a]\right)$ iff $\left([x],[\alpha^{-1}_{X}(y)]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
 iff $\left([x],\alpha^{-1}_{Y}\left([y]\right)\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}$
 iff $\left([x],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}\circ\alpha_{Y}={\sim}\varphi\left([a]\right).$
+The proof that $\varphi$ preserves -- is similar, and uses the definition of $\alpha_{Y}$.
 Theorem 4.12 finds an immediate application in this final subsection. By combining it with Definition 4.1, we get the following result:
@@
 $\mathbb{P_{U}}(\mathsf{RSM})=\mathsf{RSM}$.
-Proof 4.14
+Proof 4.15
 Consider $\{\,\mathbf{A}_{i}\mid i\in I\,\}\subseteq\mathsf{RSM}$.
@@
 $\varphi(a\cdot^{\mathbf{A}}c)(i)=\varphi_{i}\left((a\cdot^{\mathbf{A}}c)(i)\right)=\varphi_{i}\left(a(i)\cdot^{\mathbf{A}_{i}}c(i)\right)=\varphi_{i}(a(i))\cdot^{(\mathbf{B}_{i})_{r}}\varphi_{i}(c(i))=\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i)).$
-Also observe that $\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i))=\left(\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right)(i)$. Using this in the last step below, we can now see that
-$\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)=\overline{\varphi}\left([a]\right)\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\overline{\varphi}\left([c]\right)=[\varphi(a)]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}[\varphi(c)]=\left[\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right]=\left[\varphi(a\cdot^{\mathbf{A}}c)\right]$
-and hence $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)$.
+Also observe that $\varphi_{i}(a(i))\cdot^{\mathbf{B}_{i}}\varphi_{i}(c(i))=\left(\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right)(i)$.
+Using this in the last step below, we can now see that $\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)=\overline{\varphi}\left([a]\right)\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}\overline{\varphi}\left([c]\right)=[\varphi(a)]\cdot^{\mathbf{B}/\theta_{\mathcal{F}}}[\varphi(c)]=\left[\varphi(a)\cdot^{\mathbf{B}}\varphi(c)\right]=\left[\varphi(a\cdot^{\mathbf{A}}c)\right]$ and hence $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a]\right)\cdot^{(\mathbf{B}/\theta_{\mathcal{F}})_{r}}\overline{\varphi}\left([c]\right)$.
 Using the same approach, it is straightforward to show that $\overline{\varphi}$ preserves the operations $\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}$, $\vee^{\mathbf{A}/\theta_{\mathcal{F}}}$, $1^{\mathbf{A}/\theta_{\mathcal{F}}}$ and ${\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}$.
@@
 Finally, we get:
-$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)$ $=\left[\varphi(a\to^{\mathbf{A}}c)\right]$
-$=\left[{\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
+$\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)=\left[\varphi(a\to^{\mathbf{A}}c)\right]$ $=\left[{\sim}^{\mathbf{B}}\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
 $={\sim}^{\mathbf{B}/\theta_{\mathcal{F}}}\left[\left(-^{\mathbf{B}}\varphi(c)\cdot^{\mathbf{B}}\varphi(a)\right)\right]$
@@
 Every Sugihara monoid is representable. That is, $\mathsf{SM}\subseteq\mathsf{RSM}$.
-Proof 4.17
+Proof 4.18
 We need to show that $\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})=\mathsf{SM}$ is contained in $\mathsf{RSM}$.
@@
 Acknowledgements
-The authors are extremely grateful to the two anonymous referees for their very insightful and knowledgeable feedback on the first version on this paper.
+The authors are extremely grateful to the three anonymous referees for their very insightful and knowledgeable feedback on the earlier versions of this paper.
 Their comments enabled us to see that additional results could be possible, and this led to the development of Section 4.

### v4

@@
 From this it follows that all Sugihara monoids are representable.
+volume: 195issue: 1-4
 1 Introduction
@@
 See Section 2.1 for a precise definition of these algebras.
-DInFL-algebras have two order-reversing operations $\sim$ and --.
-If, for all $a\in\mathbf{A}$, ${\sim}a=-a$ then the algebra $\mathbf{A}$ is said to be cyclic.
+DInFL-algebras have two order-reversing operations $\sim$ and -.
+If, for all $\in\mathbf{A}, a=a{\sim}$a=-a then the algebra $\mathbf{A}$ is said to be cyclic.
 The use of a non-identity order isomorphism $\alpha$ allows for the construction of concrete non-cyclic DInFL-algebras (see item (ii) of Theorem 2.7).
 Our representation of Sugihara monoids is part of a more general method which can also handle non-cyclic DInFL-algebras.
-[…]
+Before giving an outline of the content of this paper, we comment on the recent representation of Sugihara monoids given by Kramer and Maddux [14].
+That paper provides an alternative solution to the problem of defining an involutive order-reversing operation on relations that has a fixed point.
+The Sugihara chains are embedded into proper relations algebras (see [14, Definition 4] for the definition of the binary relations that will be the elements of the Sugihara chain).
+For an abstract relation algebra, they define relativized subreducts [14, Definition 8] where certain operations from the original relation algebra are retained, while others are replaced by new term-definable operations.
+This notion of relativized subreduct is then implemented on algebras of relations (see [14, Theorem 2]) in order to obtain a new $\sim$} operation that will have a fixed point, hence allowing for the representation of odd Sugihara monoids.
+A difference between our approach and the one in [14, Theorem 5] is that our representations are embedded into distributive InFL-algebra of upsets (see Section 2).
+In their case, the concrete odd Sugihara chain is embedded into a proper relation algebra.
+A feature of our construction is that the relations which are the elements of the representation are all weakening relations with respect to a partial order.
 Our paper is laid out as follows.
@@
 2.1 Sugihara monoids and DInFL-algebras
-A residuated lattice is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,\backslash,/\rangle$ such that $\langle A,\wedge,\vee\rangle$ is a lattice, $\langle A,\cdot,1\rangle$ is a monoid and the following holds for all $a,b,c\in A$:
-$a\cdot b\leqslant c\quad\Longleftrightarrow\quad a\leqslant c/b\quad\Longleftrightarrow\quad b\leqslant a\backslash c.$
-[…]
-[…]
+A residuated lattice is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,\backslash$$\rangle$ such that $\langle A,\wedge,\vee\rangle$ is a lattice, $\langle A,\cdot,1\rangle$ is a monoid and the following holds for all $a,b,c\in A$:
+$\c.a\cdot b\leqslant c\quad\Longleftrightarrow\quad a\leqslant c/b\quad\Longleftrightarrow\quad b\leqslant a\backslash$ c.
+If the monoid operation of a residuated lattice $\mathbf{A}$ is commutative, then both residuals $\backslash$ and // (sometimes denoted $\to$ and $\leftarrow$) coincide.
+If $a\cdot a=a$ is satisfied for all $a\in A$, then $\mathbf{A}$ is an idempotent residuated lattice.
+A Sugihara monoid is an algebra $\mathbf{A}=\langle A,\wedge,\vee,\cdot\to,1,{\sim}\rangle$ such that $\langle A,\wedge,\vee,\cdot\to,1\rangle$ is a commutative distributive idempotent residuated lattice, and for all $a,b\in A$ the identities ${\sim}{\sim}a=a$ and $a\to{\sim}b=b\to{\sim}a$ are satisfied.
+(A commutative residuated lattice satisfying those two identities is often called an involutive commutative residuated lattice.)
+It was shown by Dunn [1] that the variety of Sugihara monoids is generated by algebras that are linearly ordered, and thus the linear ordered algebras are of particular interest.
+For our explicit representations in Section 3, we will only be considering finite Sugihara chains.
+These are algebras $\mathbf{S}_{n}$ for $n\in\omega$ and $n\geqslant 2$.
+If $n=2k$ for $k>0$ then $S_{n}=\{a_{-k},\dots,a_{-1},a_{1},\dots,a_{k}\}$, and if $n=2k+1$ for $k>0$ then $S_{n}=\{a_{-k},\dots,a_{-1},a_{0},a_{1},\dots a_{k}\}$.
+The meet and join are defined as expected: $a_{i}\wedge a_{j}=a_{\text{min}\{i,j\}}$ and $a_{i}\vee a_{j}=a_{\text{max}\{i,j\}}$.
+The unary operation $\sim$} is defined by ${\sim}a_{j}=a_{-j}$.
+The monoid operation can be defined as:
 $a_{i}\cdot a_{j}=\begin{cases}a_{i}&\text{if }\,|j|<|i|\\ a_{j}&\text{if }\,|i|<|j|\\ a_{\text{min}\{i,j\}}&\text{if }\,|j|=|i|.\end{cases}$
 Lastly, the implication $\to$ is defined by
-$a_{i}\to a_{j}=\begin{cases}{\sim}a_{i}\vee a_{j}&\text{if }i\leqslant j\\ {\sim}a_{i}\wedge a_{j}&\text{if }i>j.\end{cases}$
+[…]
 If $n$ is odd then the monoid identity is given by $1=a_{0}$, and if $n$ is even, then the monoid identity is $1=a_{1}$.
@@
 (When we talk of a Sugihara chain, we mean a linearly ordered Sugihara monoid, using slightly different terminology from that used in [16] and [14] where a Sugihara chain is a linearly ordered Sugihara algebra.)
-[…]
-[…]
+The concrete algebras we describe in Section 2.3 will be special residuated lattices, which we describe briefly here.
+A Full Lambek algbebra (cf.
+[11]) is a residuated lattice $\mathbf{L}$ that has an additional constant 0 (about which nothing is assumed) in the signature.
+We use the common abbreviation FL-algebra.
+Using the residuals, two unary operations $\sim$} and - can be defined.
+These are known as linear negations and are defined for $a\in L$ by ${\sim}a=a\backslash 0$ and a=0/a-a=0/a.
+If $\sim$} and - satisfy the involutive property, i.e.
+${\sim}a=a={\sim}{-}a$ for all $a\in L$, then $\langle L,\wedge,\vee,\cdot,1,0,\backslash$$\rangle$ is an InFL-algebra.
+When the underlying lattice is distributive, we use the abbreviation DInFL-algebra.
+If ${\sim}a=-a$ for all $a\in L$ then it is a cyclic InFL-algebra.
+We prefer to use the linear negations rather than the residuals in the signature as it is known that such an equivalent formulation is available [8, Lemma 2.2].
+In addition, that equivalent formulation allows us to leave the 0 out of the signature.
+That is, we will consider DInFL-algebras as algebras of the following signature: $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,{\sim},$-$\rangle$.
+Notice that if a Sugihara monoid uses an alternative signature with $\sim$} then it will be a DInFL-alegbra.
+Moreover, it will be a commutative cyclic idempotent DInFL-algebra.
 2.2 Binary relations
@@
 2.3 Constructing InFL-algebras of relations
-[…]
+In previous work [3, Section 3], we used the construction outlined below to produce concrete algebras of binary relations which formed distributive quasi relation algebras (DqRAs).
+These algebras were described by Galatos and Jipsen [8] and are algebras of the form $\mathbf{A}=\langle A,\wedge,\vee,\cdot,1,0,{\sim},-,^{\prime}\rangle$.
+The unary operations $\sim$},- and are dual lattice isomorphisms satisfying some additional conditions.
+We do not go into further detail here about the definition of quasi relation algebras as we will not make use of the full signature of these algebras.
+The construction outlined below will produce distributive involutive FL-algebras.
 The definitions and results below all involve an equivalence relation $E$ on a set $X$.
@@
 Notice that when $E=X^{2}$ then $\langle X^{2},\preccurlyeq\rangle$ is exactly the product poset $\mathbf{X}^{\partial}\times\mathbf{X}$ where the order on $\mathbf{X}^{\partial}$ is $\geqslant$.
-[…]
+We note that when $E=X^{2}$ the lattice constructed above is identical to that used by Galatos and Jipsen [9, 10], and more recently by Jipsen and Šemrl [13].
+In order to represent odd Sugihara monoids as algebras of relations, we will make use of an order automorphism $\alpha$ on the poset $\mathbf{X}$.
+It is this $\alpha$ that will allow us to define a unary operation $\sim$} on $\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ and an identity $1\in\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ such that ${\sim}1=1$.
 The lemma below is stated in the original source [3] with an additional map $\beta:X\to X$ which is a dual order automorphism.
@@
 […]
-The next lemma shows that if $\alpha$ is an order automorphism of a poset $\langle X,\leqslant\rangle$ with an equivalence relation $E$ such that ${\leqslant}\subseteq E$, then $\alpha$ can be used to define an element of $\mathsf{Up}(\mathbf{E})$ to act as a 00 that will give rise to an InFL-algebra.
+The next lemma shows that if $\alpha$ is an order automorphism of a poset $\langle X,\leqslant\rangle$ with an equivalence relation $E$ such that ${\leqslant}\subseteq E$, then $\alpha$ can be used to define an element of $\mathsf{Up}(\mathbf{E})$ to act as a 0 that will give rise to an InFL-algebra.
 Lemma 2.5 ([3, Lemma 3.9])
@@
 Then $\alpha\,\circ\leqslant^{c\smile}=$ $\leqslant^{c\smile}\circ\,\alpha\in\mathsf{Up}\left(\mathbf{E}\right)$.
-We will now add $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ to the signature of the distributive residuated lattice $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,\backslash,/\right\rangle$ to obtain the FL-algebra $\left\langle\mathsf{Up}\left(\mathbf{E}\right),\cap,\cup,\circ,\leqslant,0,\backslash,/\right\rangle$.
-The next lemma gives an alternative way of calculating ${\sim}R$ and $-R$ if we set $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$.
+[…]
 Lemma 2.6 ([3, Lemma 3.10])
 Let $\mathbf{X}=\langle X,\leqslant\rangle$ be a poset with $E$ an equivalence relation and $\alpha:X\rightarrow X$ an order automorphism of $\mathbf{X}$ where ${\leqslant}\subseteq E$ and $\alpha\subseteq E$.
-If $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ then ${\sim}R=R^{c\smile}\circ\alpha$ and $-R=\alpha\circ R^{c\smile}$ for all $R\in\mathsf{Up}\left(\langle X^{2},\preccurlyeq\rangle\right)$.
-[…]
+If $0=\alpha\,\circ\leqslant^{c\smile}\,=\,\leqslant^{c\smile}\circ\,\alpha$ then ${\sim}R=R^{c\smile}\circ\alpha$ and $R=\alpha\circ R^{c\smile}$ for all $R\in\mathsf{Up}\left(\langle X^{2},\preccurlyeq\rangle\right)$.
+Using the alternative method of calculating $\sim$} and {-} from Lemma 2.6, greatly simplifies the proof of the main theorem of this background section.
 Theorem 2.7 ([3, Theorem 3.12])
@@
 Example 2.10
-[…]
+Consider the three-element Sugihara chain $\mathbf{S}_{3}$.
+The underlying set is given by $S_{3}=\{a_{-1},a_{0},a_{1}\}$ where $a_{0}$ is the identity of the monoid operation.
+Consider the two-element poset $X=\left\{x,y\right\}$ with order $\leqslant\,=\mathrm{id}_{X}$.
+We let $E=X^{2}$ and define the order automorphism $\alpha=\left\{\left(x,y\right),\left(y,x\right)\right\}$.
+We get that $\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ has as its underlying lattice the 16-element Boolean lattice.
+It is straightforward to calculate that ${\leqslant}={\sim}{\leqslant}={\leqslant}^{c\smile}\circ\,\alpha$ and then to show that the set of binary relations $\{\varnothing,\leqslant,X^{2}\}$ is closed under $\cap$, $\cup$, $\circ$, $\Rightarrow$ and $\sim$}.
+Define a map $h:S_{3}\to\mathsf{Up}(\langle X^{2},\preccurlyeq\rangle)$ by $h(a_{-1})=\varnothing$, $h(a_{0})={\leqslant}$ and $h(a_{1})=X^{2}$.
+It is easy to check that this is indeed an embedding of $\mathbf{S}_{3}$ into the direct reduct of $\mathfrak{D}(\mathbf{E})$.
 The next example is based on the representation of the finite even Sugihara chains $\mathbf{S}_{2n+2}$ ($n\in\mathbb{Z}^{+}$) by Maddux [16, Theorem 6.2].
@@
 Now define binary relations $\mathrm{id}_{\mathbb{Q}^{n}}$ and $L_{1}$ on $\mathbb{Q}^{n}$ as follows:
-[…]
-[…]
+$\left(p,q\right)\in\mathrm{id}_{\mathbb{Q}^{n}}\qquad$ $\textnormal{ iff }\qquad p=q$
+$(p,q)\in L_{1}\qquad$ $\textnormal{ iff }\qquad p_{1}<q_{1}.$
 For $i\in\{2,\ldots,n\}$, the relation $L_{i}$ on $\mathbb{Q}^{n}$ is defined follows:
@@
 $T_{i}=\begin{cases}\varnothing&\text{if }\,i=-n-1\\ \displaystyle\bigcup_{j=1}^{n+1+i}L_{j}&\text{if }\,-n\leqslant i\leqslant-1\\ \leqslant_{n}&\text{if }\,i=1\\ \left(T_{-i}\right)^{c\smile}&\text{if }\,2\leqslant i\leqslant n+1\end{cases}$
-These relations are upsets of $\langle E,\preccurlyeq\rangle$ and form a Sugihara chain of length $22n+$ 2, i.e.,
+These relations are upsets of $\langle E,\preccurlyeq\rangle$ and form a Sugihara chain of length $2n+2$, i.e.,
 $\varnothing=T_{-n-1}\subseteq T_{-n}\subseteq\cdots\subseteq T_{-1}\subseteq T_{1}\subseteq\cdots\subseteq T_{n}\subseteq T_{n+1}=\left(\mathbb{Q}^{n}\right)^{2}.$
@@
 Figure 1: The poset $\langle X,\leqslant_{X}\rangle$ from a chain $\langle C,\leqslant\rangle$ with order automorphism $\alpha:X\to X$.
-[…]
+The first lemma in this section establishes that when we have a poset $\langle X,\leqslant_{X}\rangle$ constructed from a chain as above, the order $\leqslant_{X}$ will be fixed by the $\sim$} operation described in Theorem 2.7.
+Note that $(x^{b},y^{d})\in{\sim}R$ if and only if $(y^{-d},x^{b})\notin R$.
+In particular, $(x^{b},y^{d})\in 0$ if and only if $y^{-d}\nleqslant x^{b}$.
 Lemma 3.1
@@
 Then we have $y^{-d}\not\leqslant x^{b}$.
-If $b=-d$, then $y\not\leqslant x$, so $x<y$.
+If b=db=-d, then $y\not\leqslant x$, so $x<y$.
 Hence, $x^{b}\leqslant_{X}y^{d}$.
@@
 Define binary relations $\mathrm{id}_{\mathbb{Q}^{n}}$ and $L_{1}$ on $\mathbb{Q}^{n}$ by
-[…]
-[…]
+$\left(p,q\right)\in\mathrm{id}_{\mathbb{Q}^{n}}$  $\quad\textnormal{ iff }\quad p=q$
+$\left(p,q\right)\in L_{1}$  $\quad\textnormal{ iff }\quad p_{1}<q_{1}$
 and for $i\in\left\{2,\ldots,n\right\}$, define binary relations $L_{i}$ on $\mathbb{Q}^{n}$ by
-[…]
+$\left(p,q\right)\in L_{i}$  $\quad\textnormal{ iff }\quad\left(p_{1},\ldots,p_{i-1}\right)=\left(q_{1},\ldots,q_{i-1}\right)\textnormal{ and }p_{i}<q_{i}.$
 Now set $\leqslant_{n}:=\bigcup\left\{\mathrm{id}_{\mathbb{Q}^{n}},L_{1},\ldots,L_{n}\right\}$.
@@
 The fact that $\leqslant_{X}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ follows from Lemma 2.4.
-It would also follow from Lemmas 2.3 and 2.6 that $R_{i}$ is an upset for $i\in\left\{1,\ldots,n\right\}$, if we can show that $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ when $-n\leqslant i<0$.
+It would also follow from Lemmas 2.3 and 2.6 that $R_{i}$ is an upset for $i\in\left\{1,\ldots,n\right\}$, if we can show that $R_{i}$ is an upset of $\left\langle X^{2},\preccurlyeq\right\rangle$ when $n\leqslant i<0$.
 Let $(p^{b},q^{d})\in R_{i}$ and assume $\left(p^{b},q^{d}\right)\preccurlyeq\left(r^{e},s^{f}\right)$.
@@
 Hence, $\left(r,s\right)\in L_{\text{min}\{j,k,\ell\}}$, which means $\left(r^{e},s^{f}\right)\in U_{\text{min}\{j,k,\ell\}}\subseteq R_{i}$.
-Notice that when defining $R_{i}$ for $i\in\left\{1,\ldots,n+1\right\}$ we chose to use the unary order-reversing operation $\sim$ instead of --.
+Notice that when defining $R_{i}$ for $i\in\left\{1,\ldots,n+1\right\}$ we chose to use the unary order-reversing operation $\sim$ instead of -.
 Since $\alpha\neq\text{id}_{X}$ we know from Theorem 2.7 that the algebra will not be cyclic and hence ${\sim}R\neq-R$ in general.
@@
 Lemma 3.6
-For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $-R_{i}={\sim}R_{i}$.
+For each $i\in\left\{-n-1,\ldots,n+1\right\}$, we have $R_{i}={\sim}R_{i}$.
 Proof 3.7
 If $i=0$, the claim follows from Lemma 2.5. Now consider the following cases:
-Case 1: $-n-1\leqslant i<0$.
+Case 1: $n-1\leqslant i<0$.
 Let $\left(p^{b},q^{d}\right)\notin-R_{i}$.
@@
 The reverse inclusion can be proved in a similar way.
-Case 2: $10<i\leqslant n+$ 1. This case follows from the previous case and the fact that -- and $\sim$ satisfy the involutive property.
+Case 2: $0<i\leqslant n+1$. This case follows from the previous case and the fact that - and $\sim$ satisfy the involutive property.
 We want our set of relations to have the structure of a Sugihara chain, therefore we need to confirm that they form a linear order when ordered by inclusion.
@@
 Proof 3.11
-This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for $i=-n-1$. We now consider the following two cases:
-Case 1: $-n\leqslant i<0$.
+This is clearly true for $i=0$ and $i=n+1$. Since $\varnothing$ is vacuously transitive, the statement also holds for i=n1i=-n-1. We now consider the following two cases:
+Case 1: $n\leqslant i<0$.
 Let $\left(p^{b},r^{e}\right)\in R_{i}$ and $\left(r^{e},q^{d}\right)\in R_{i}$.
@@
 Proof 3.13
-If $i=-n-1$ (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
-Case 1: $-n-1<i<j<0$.
+If i=n1i=-n-1 (i.e. $R_{i}=\varnothing$) then we clearly have $R_{i}\circ R_{j}=R_{i}$. The case for $j=0$ follows from Lemma 2.2. So we consider the following cases:
+Case 1: n1<i<j<0-n-1<i<j<0.
 Since $R_{j}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{i}$.
@@
 This gives us $(p^{b},q^{d})\in R_{i}\circ R_{j}$ and so $R_{i}\circ R_{j}=R_{i}$.
-Case 2: $-n-1<i<0$, $j>0$.
+Case 2: n1<i<0-n-1<i<0, $j>0$.
 Since ${\leqslant_{X}}\subseteq R_{j}$, we get $R_{i}\subseteq R_{i}\circ R_{j}$.
@@
 Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}$.
-Case 3: $i>0$, $-n-1<j<0$.
+Case 3: $i>0$, n1<j<0-n-1<j<0.
 Since $R_{j}\subseteq{\leqslant_{X}}$ we get $R_{i}\circ R_{j}\subseteq R_{i}$.
@@
     For all $k\in\left\{2,\ldots,n+1-i\right\}$, we have $p_{k}=q_{k}$.
-Let $r=\left(q_{1},\ldots,q_{n+1-i},q_{n+2-i}-1,\ldots,q_{n}-1\right)$.
+Let r = (q_1, …, q_n+1-i, q_n+2-i-1,…, q_n-1).
 Then $(r,q)\in L_{n+2-i}$, and so $\left(r^{b},q^{d}\right)\in U_{n+2-i}\subseteq\bigcup_{k=1}^{n+1+j}\subseteq R_{j}$.
@@
 Proof 3.15
-If $j=-n-1$, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
-Case 1: $-n-1<j<i<0$.
+If j=n1j=-n-1, then $R_{j}=\varnothing$, and so $R_{i}\circ R_{j}=R_{j}$. If $i=0$, we have $R_{i}=\;\leqslant_{X}$, so $R_{i}\circ R_{j}=R_{j}$ by Lemma 2.4. We now consider the following cases:
+Case 1: n1<j<i<0-n-1<j<i<0.
 Since $R_{i}\subseteq{\leqslant_{X}}$, we get $R_{i}\circ R_{j}\subseteq R_{j}$.
@@
 Setting $r=\left(q_{1},\ldots,q_{k-1},r_{k},0,\ldots,0\right)$, we obtain $\left(p^{b},r^{b}\right)\in U_{k}\subseteq R_{j}\subseteq R_{i}$ and $\left(r^{b},q^{d}\right)\in U_{k}\subseteq R_{j}$, which means $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}$.
-Case 2: $i>0,-n-1<j<0$.
+Case 2: i>0,n1<j<0i>0,-n-1<j<0.
 Since ${\leqslant_{X}}\subseteq R_{i}$, we get $R_{j}\subseteq R_{i}\circ R_{j}$.
@@
   2. (b)
-    For all $k\in\left\{1,\ldots,n+1-j\right\}$, we have $p_{k}=q_{k}$. Let
-    $r=\left(p_{1},\ldots,p_{n+1-j},p_{n+2-j}+1,\ldots,p_{n}+1\right).$
+    For all $k\in\left\{1,\ldots,n+1-j\right\}$, we have $p_{k}=q_{k}$.
+Let r = (p_1, …, p_n+1-j, p_n+2-j+1, …, p_n + 1).
 Then $\left(p,r\right)\in L_{n+2-j}$, and so $\left(p^{b},r^{b}\right)\in U_{n+2-j}\subseteq\bigcup_{k=1}^{n+1+i}U_{k}=R_{i}$.
@@
 Proof 3.17
-The claim clearly follows for the cases $i=j=0$ and $i=j=-n-1$. So consider the following cases:
-Case 1: $-n-1<i=j<0$.
+The claim clearly follows for the cases $i=j=0$ and i=j=n1i=j=-n-1. So consider the following cases:
+Case 1: n1<i=j<0-n-1<i=j<0.
 Let $\left(p^{b},q^{d}\right)\in R_{i}\circ R_{j}=R_{i}\circ R_{i}$.
@@
 This gives us $(p^{b},q^{d})\in R_{i}\circ R_{j}$.
-Case 2: $-n-1<i<0$, $j>0$.
+Case 2: n1<i<0-n-1<i<0, $j>0$.
 Since $\leqslant_{X}\;\subseteq R_{j}$, we have $R_{\text{min}\{i,j\}}=R_{i}\subseteq R_{i}\circ R_{j}$.
@@
 Hence, $(p,q)\in L_{m}$, and so $\left(p^{b},q^{d}\right)\in U_{m}\subseteq R_{i}=R_{\text{min}\{i,j\}}$.
-Case 3: $-n-1<j<0,i>0$.
+Case 3: n1<j<0,i>0-n-1<j<0,i>0.
 Since $\leqslant_{X}\;\subseteq R_{i}$, we have $R_{\text{min}\{i,j\}}=R_{j}\subseteq R_{i}\circ R_{j}$.
@@
 $\mathbb{S}_{2n+3}:=\langle\mathcal{R}_{2n+3},\cap,\cup,\circ,\Rightarrow,R_{0},{\sim}\rangle$
-[…]
+where $\Rightarrow$ is defined, as in the abstract case, in terms of $\cap$, $\cup$ and $\sim$}.
 Theorem 3.19
@@
 Proof 3.20
-[…]
+The fact that the ordered set forms a chain is Lemma 3.8.
+Comparing Proposition 3.18 with the definition of $\mathbf{S}_{2n+3}$ in Section 2.1 shows that the monoid operation is defined as it should be for an odd Sugihara chain.
+The fact that $\sim$} is order-reversing follows from the fact that $R_{i}={\sim}R_{-i}$ and Lemma 3.8.
+The isomorphism is given by $\varphi:S_{2n+3}\to\mathcal{R}_{2n+3}$ where $\varphi(a_{i})=R_{i}$.
 4 Representing all Sugihara monoids via weakening relations
@@
 Now consider the set of equivalence classes $Y=X/E_{\mathcal{F}}$.
-Define binary relations $\leqslant_{Y}$ and $E_{Y}$ on $Y$ by setting, for all $[x],[y]\in Y$,
+Define binary relations $\leqslant_{Y}$ and $E_{Y}$ on $Y$ by setting, for all $x],[y]\in Y$,
 $[x]\leqslant_{Y}[y]\qquad$ $\textnormal{iff}\qquad\left\{i\in I\mid x(i)\leqslant_{i}y(i)\right\}\in\mathcal{F},$
@@
 Suppose further that for each $i\in I$, the map $\alpha_{i}:X_{i}\to X_{i}$ is an order automorphism of $\mathbf{X}_{i}$ such that $\alpha_{i}\subseteq E_{i}$.
-Define $\alpha_{Y}:Y\to Y$ by setting, for each $[x]\in Y$,
+Define $\alpha_{Y}:Y\to Y$ by setting, for each $x]\in Y$,
 $\alpha_{Y}\left([x]\right)=\left[\alpha_{X}(x)\right]$
@@
   4. (iv)
-    If $0=\alpha_{Y}\circ\left(\leqslant_{Y}\right)^{c\smile}=\left(\leqslant_{Y}\right)^{c\smile}\circ\alpha_{Y}$, then $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},0,\sim,-\right\rangle$ is a distributive InFL-algebra.
+[…]
 Proof 4.6
@@
 We may assume that $X_{i}\neq\varnothing$ for all $i\in I$ since the direct product of a family of algebras from $\mathsf{EDInFL}$ is isomorphic to another such product from which all factors of the form $\mathfrak{D}\left(\left\langle\varnothing,\varnothing\right\rangle\right)$ have been deleted.
-Now, for each $[a]\in A/\theta_{\mathcal{F}}$, define
+Now, for each $a]\in A/\theta_{\mathcal{F}}$, define
 $\left([x],[y]\right)\in R_{[a]}\qquad\textnormal{iff}\qquad\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}.$
 Lemma 4.8
-For any $[a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
+For any $a]\in A/\theta_{\mathcal{F}}$, the set $R_{[a]}$ is an upset of $\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle$.
 Proof 4.9
@@
 The first statement implies that we have $\left\{i\in I\mid\left(x(i),y(i)\right)\in\varphi_{i}(a(i))\right\}\in\mathcal{F}$.
-From the second part we get $[u]\leqslant_{Y}[x]$ and $[y]\leqslant_{Y}[v]$, which means $\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}$.
+From the second part we get $u]\leqslant_{Y}[x]$ and $y]\leqslant_{Y}[v]$, which means $\left\{i\in I\mid u(i)\leqslant_{i}x(i)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid y(i)\leqslant_{i}v(i)\right\}\in\mathcal{F}$.
 Since $\mathcal{F}$ is closed under taking intersections, we have
@@
 By Lemma 4.8 we know that $\varphi$ is well-defined.
-Now assume that $[a]\neq[b]$.
+Now assume that $a]\neq[b]$.
 This implies $(a,b)\notin\theta_{\mathcal{F}}$ and hence $\{i\in I\mid a(i)=b(i)\}\notin\mathcal{F}$.
@@
 Proof 4.13
-We have to show that the ultraproduct $\mathbf{A}/\theta_{\mathcal{F}}$ of $\left\{\mathbf{A}_{i}\mid i\in I\right\}$ is representable.
-Recall that each of the $\mathbf{A}_{i}$ is embedded via $\varphi_{i}$ into an algebra of the form $\mathfrak{D}\left(\langle E_{i},\preccurlyeq_{i}\rangle\right)$.
-Each $\langle E_{i},\preccurlyeq_{i}\rangle$ comes from a poset $\mathbf{X}_{i}$ equipped with the necessary $E_{i}$ and $\alpha_{i}$.
-The definitions after Theorem 4.3 describe the structure $\langle E_{Y},\preccurlyeq_{Y}\rangle$, and, by Theorem 4.5(iv), $\left\langle\mathsf{Up}\left(\left\langle E_{Y},\preccurlyeq_{Y}\right\rangle\right),\cap,\cup,\circ,\leqslant_{Y},\sim,-\right\rangle$ is a distributive InFL-algebra.
+[…]
 From Lemma 4.10 we have that $\varphi$ is injective.
@@
 To show that $\varphi$ preserves the monoid operation, first let $\left([x],[y]\right)\in\varphi\left([a]\right)\circ\varphi\left([b]\right)$.
-Then there is some $[z]\in Y$ such that $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$.
+Then there is some $z]\in Y$ such that $\left([x],[z]\right)\in\varphi\left([a]\right)$ and $\left([z],[y]\right)\in\varphi\left([b]\right)$.
 Hence, we obtain $\left\{i\in I\mid\left(x(i),z(i)\right)\in\varphi_{i}\left(a(i)\right)\right\}\in\mathcal{F}$ and $\left\{i\in I\mid\left(z(i),y(i)\right)\in\varphi_{i}\left(b(i)\right)\right\}\in\mathcal{F}$, which means
@@
 iff $\left([x],[y]\right)\in\left(\varphi\left([a]\right)\right)^{c\smile}\circ\alpha_{Y}={\sim}\varphi\left([a]\right).$
-The proof that $\varphi$ preserves -- is similar, and uses the definition of $\alpha_{Y}$.
+The proof that $\varphi$ preserves - is similar, and uses the definition of $\alpha_{Y}$.
 Theorem 4.12 finds an immediate application in this final subsection. By combining it with Definition 4.1, we get the following result:
@@
 We show that $\overline{\varphi}$ is bijective.
-For $[b]\in B/\theta_{\mathcal{F}}$, since $b\in B$, there exists $a\in A$ with $\varphi(a)=b$.
+For $b]\in B/\theta_{\mathcal{F}}$, since $b\in B$, there exists $a\in A$ with $\varphi(a)=b$.
 So $\overline{\varphi}\left([a]\right)=\left[\varphi(a)\right]=[b]$.
-If $a,c\in A$ with $[a]\neq[c]$, then $(a,c)\notin\theta_{\mathcal{F}}$.
+If $a,c\in A$ with $a]\neq[c]$, then $(a,c)\notin\theta_{\mathcal{F}}$.
 Thus $\{\,i\in I\mid a(i)=c(i)\,\}\notin\mathcal{F}$, so $\{\,i\in I\mid a(i)\neq c(i)\,\}\in\mathcal{F}$.
@@
 Hence $\{\,i\in I\mid a(i)\neq c(i)\,\}\subseteq\{\,i\in I\mid\varphi(a)(i)\neq\varphi(c)(i)\,\}$ so the latter set is in $\mathcal{F}$.
-Now $\{\,i\in I\mid\varphi(a)(i)=\varphi(c)(i)\,\}\notin\mathcal{F}$, so $(\varphi(a),\varphi(c))\notin\theta_{\mathcal{F}}$ and therefore $[\varphi(a)]\neq[\varphi(c)]$, i.e.
+Now $\{\,i\in I\mid\varphi(a)(i)=\varphi(c)(i)\,\}\notin\mathcal{F}$, so $(\varphi(a),\varphi(c))\notin\theta_{\mathcal{F}}$ and therefore $\varphi(a)]\neq[\varphi(c)]$, i.e.
 $\overline{\varphi}\left([a]\right)\neq\overline{\varphi}\left([c]\right)$.
-Next, we will show that $\overline{\varphi}$ preserves the fusion operation. For $[a],[c]\in A/\theta_{\mathcal{F}}$, we have
+Next, we will show that $\overline{\varphi}$ preserves the fusion operation. For $a],[c]\in A/\theta_{\mathcal{F}}$, we have
 $\overline{\varphi}\left([a]\cdot^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left(\left[a\cdot^{\mathbf{A}}c\right]\right)=\left[\varphi\left(a\cdot^{\mathbf{A}}c\right)\right].$
@@
 Using the same approach, it is straightforward to show that $\overline{\varphi}$ preserves the operations $\wedge^{\mathbf{A}/\theta_{\mathcal{F}}}$, $\vee^{\mathbf{A}/\theta_{\mathcal{F}}}$, $1^{\mathbf{A}/\theta_{\mathcal{F}}}$ and ${\sim}^{\mathbf{A}/\theta_{\mathcal{F}}}$.
-Now consider the operation $\to^{\mathbf{A}/\theta_{\mathcal{F}}}$ on $\mathbf{A}/\theta_{\mathcal{F}}$. For $[a]$, $[c]\in A/\theta_{\mathcal{F}}$, we have
+Now consider the operation $\to^{\mathbf{A}/\theta_{\mathcal{F}}}$ on $\mathbf{A}/\theta_{\mathcal{F}}$. For $[a]$, $c]\in A/\theta_{\mathcal{F}}$, we have
 $\overline{\varphi}\left([a]\to^{\mathbf{A}/\theta_{\mathcal{F}}}[c]\right)=\overline{\varphi}\left([a\to^{\mathbf{A}}c]\right)=\left[\varphi(a\to^{\mathbf{A}}c)\right].$
@@
 Every algebra can be embedded in an ultraproduct of its finitely generated subalgebras.
-The finitely generated subalgebras of $\mathbf{S}=(\mathbb{Z},\wedge,\vee,\cdot,\to,1,{\sim})$ are the algebras $\mathbf{S}_{2n+1}$ for $n\in\omega$.
-For the infinite Sugihara monoid $\mathbf{S}^{*}=(\mathbb{Z}{\setminus}\{0\},\wedge,\vee,\cdot,\to,1,{\sim})$, its finitely generated subalgebras are $\mathbf{S}_{2n+2}$ for $n\in\omega$.
+The finitely generated subalgebras of $\mathbf{S}=\langle\mathbb{Z},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$ are the algebras $\mathbf{S}_{2n+1}$ for $n\in\omega$.
+For the infinite Sugihara monoid $\mathbf{S}^{*}=\langle\mathbb{Z}{\setminus}\{0\},\wedge,\vee,\cdot,\to,1,{\sim}\rangle$, its finitely generated subalgebras are $\mathbf{S}_{2n+2}$ for $n\in\omega$.
 Lastly, recall that the quasivariety of Sugihara monoids is generated by $\{\mathbf{S},\mathbf{S}^{*}\}$ (cf.
@@
 (The trivial algebra $\mathbf{S}_{1}$ is of course representable as it is a subalgebra of $\mathbf{S}_{3}$.)
-[…]
-[…]
+Let $I=\{\,2k+1\mid k\in\omega\,\}$ and $\mathbf{A}=\prod\{\,\mathbf{S}_{i}\mid i\in I\,\}$.
+By Theorem 4.16, there exists an ultrafilter $\mathcal{F}$ on $I$ and an embedding $e_{1}:\mathbf{S}\hookrightarrow\mathbf{A}/\theta_{\mathcal{F}}$.
+Since $\mathbf{A}/\theta_{\mathcal{F}}\in\mathbb{P_{U}}(\mathsf{RSM})$ we get $\mathbf{S}\in\mathbb{ISP_{U}}(\mathsf{RSM})\subseteq\mathsf{RSM}$.
+Next, let $J=\{\,2n+2\mid n\in\omega\,\}$ and $\mathbf{B}=\prod\{\,\mathbf{S}_{j}\mid j\in J\,\}$.
+By Theorem 4.16, there is an ultrafilter $\mathcal{G}$ on $J$ and an embedding $e_{2}:\mathbf{S}^{*}\hookrightarrow\mathbf{B}/\theta_{\mathcal{G}}$.
+Again, since $\mathbf{B}/\theta_{\mathcal{G}}\in\mathbb{P_{U}}(\mathsf{RSM})$ we get $\mathbf{S}^{*}\in\mathbb{ISP_{U}}(\mathsf{RSM})\subseteq\mathsf{RSM}$.
 Since $\mathsf{SM}=\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})$ and $\{\mathbf{S},\mathbf{S}^{*}\}\subseteq\mathsf{RSM}$ we get $\mathsf{SM}=\mathbb{ISP}(\mathbf{S},\mathbf{S}^{*})\subseteq\mathbb{ISP}(\mathsf{RSM})=\mathsf{RSM}$.